### v1

# Leverage Knowledge Graph and Large Language Model for Law Article Recommendation: A Case Study of Chinese Criminal Law

Abstract
Court efficiency is vital for social stability. However, in most countries around the world, the grassroots courts face case backlogs, with decisions relying heavily on judicial personnel’s cognitive labor, lacking intelligent tools to improve efficiency. To address this issue, we propose an efficient law article recommendation approach utilizing a Knowledge Graph (KG) and a Large Language Model (LLM). Firstly, we propose a Case-Enhanced Law Article Knowledge Graph (CLAKG) as a database to store current law statutes, historical case information, and correspondence between law articles and historical cases. Additionally, we introduce an automated CLAKG construction method based on LLM. On this basis, we propose a closed-loop law article recommendation method. Finally, through a series of experiments using judgment documents from the website ”China Judgements Online”, we have improved the accuracy of law article recommendation in cases from 0.549 to 0.694, demonstrating that our proposed method significantly outperforms baseline approaches.
1 Introduction
Figure 1: The pipeline of CLAKG. The Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG).
In a modern rule-of-law society, the efficiency of court judgments is crucial for ensuring economy growth and social stability [39, 53, 41]. Meanwhile, [60] revealed that the current backlog of cases in courts is enormous, which means the workload of judicial personnel is heavy. [59] showed that there is a great need for intelligent tools to assist judicial personnel in improving the efficiency of case adjudication, since intelligent tools boost judicial efficiency by categorizing similar cases, speeding up research, and staying updated with new laws, ensuring faster and more consistent decisions.
Law article recommendation entails predicting the relevant law articles applicable to a case based on its factual description. With the advancement of computer science, many data science techniques have been applied to this task, greatly improving efficiency. Classic law article recommendation techniques, such as LSTM [44], Text-CNN [46], GRU [36], RCNN [48], HAN [62], and DPCNN [45], are based on traditional machine learning methods. These methods focus solely on the correspondence between case facts and law article ids, while overlooking the semantic information in law articles. Moreover, these methods are highly susceptible to the effects of insufficient data and imbalanced data labels. Therefore, there is a need to solve above problems.
Some researchers utilized Large Language Model (LLM) to complete the task of law article recommendation [61, 50, 56]. When engaging with extensive LLMs, users are enabled to not only retrieve the numeral and textual format of law article but also to investigate the rationale behind the applicability of these statutes to the specific case by posing supplemental inquiries. Moreover, users have the capability to probe into the rationale for the non-selection of other analogous provisions. However, [37] found that directly using LLMs is prone to hallucinations, generating incorrect law article information. Therefore, it is necessary to introduce a fully accurate legal information and case knowledge base for auxiliary analysis.
To mitigate the issue of hallucination, researchers introduced Retrieval-Augmented Generation (RAG) technology to enhance the generation capability of LLMs, by retrieving relevant information from external knowledge bases[49, 34]. Currently, RAG technology has been applied in fields such as code generation [43], autonomous driving [38], and enterprise management[51]. However, research on RAG methods in the domain of law article recommendation remains relatively underexplored. RAG is divided into two parts: retrieval and generation. In academic discussions, greater attention is often given to the retrieval component of RAG, whereas the generation process is predominantly handled by LLMs. Classic retrieval methods include TFIDF-based retrieval [32], BM25-based retrieval [54], and retrieval methods based on frozen BERT [33]. However, the above methods mainly focus on the word level and rarely use macro semantics for matching.
To deal with that problem, [35] proposed that the Knowledge Graph (KG), an unstructured database that uses a directed graph structure to store data, is highly suitable to serve as a database for RAG technology. [52] have pointed out that the integration of KGs with LLMs shows great potential for future research. [40] have utilized LLMs to automatically construct knowledge graphs, which are then used to enhance the generation capabilities of LLMs. This method performs well in terms of comprehensiveness and diversity of answers. However, the KGs generated using this method often have complex structures and lack a fixed schema, making them difficult for users to modify, such as adding new cases or removing outdated ones.
To address the above issues, we introduce knowledge graph technology and large language model technology to improve the output accuracy in law article recommendation tasks. Initially, we pre-design the schema of a Case-Enhanced Law Article Knowledge Graph (CLAKG) capable of simultaneously storing current law articles and historical cases information as a database. On this basis, we also introduce an automated CLAKG construction method based on LLM. Furthermore, we design a closed-loop human-machine collaboration method to complete the law article recommendation task. To demonstrate the effectiveness of our proposed method, we constructed a dataset of Chinese criminal law judgments. We compared the method introduced in this paper with approaches based on BERT, methods utilizing LLM, and the LLM approach based on TFIDF.
The remainder of this paper is organized as follows. Section 2 proposes a closed-loop law article recommendation framework. Section 3 utilizes the method from the second section to construct the CLAKG and compares the accuracy of this method with other baseline approaches. Lastly, the conclusions and future work are highlighted in Section 4.
2 Method
The pipeline of CLAKG-based law article recommendation is shown in Figure 1. We first establish a knowledge graph based on the designed schema. Initially, nodes and relationships are extracted from law articles and judgments in cases to construct the Law Article Knowledge Graph (LAKG) and the Adjudicated Cases Knowledge Graph (ACKG). Subsequently, by utilizing a Large Language Model (LLM), these two graphs are integrated to form CLAKG. This part is introduced in Section 2.1. Building upon this, we have developed a human-machine Collaborative law article recommendation framework. Users describe new case information, and the LLM would provide potentially applicable law articles based on keyword matching and graph embedding techniques. Moreover, the system can provide similar historical case information from the CLAKG, followed by providing most relevant law articles that match the new case using LLM. This part is introduced in Section 2.2. Following the above process, the user reviews the law articles provided by the system and corrects any inaccuracies. The new case information is used to update the CLAKG.
2.1 Case-Enhanced Law Article Knowledge Graph Construction
(a) Schema of CLAKG. Ellipses of different colors represent
different entity types.
(b) Example of CLAKG.
Figure 2: Overall schema and instantiation example.
Table 1: Entity types in Case-Enhanced Law Article Knowledge Graph
Entity Type Description
Original article of law An original article from the field of law, such as criminal law or traffic law, etc
Key information Key information in law article extracted by large language model
Law article id The article number in the law
The name of a case Case name on the judgment documents
Time of the court session The time when a court makes a decision in a case
Prosecution reason Prosecution reason on the judgment documents
The specifics of the case Case information extracted using a large language model
Table 2: Relationship types in Case-Enhanced Law Article Knowledge Graph
Relation Type Description
Key A law article contains several pieces of key information
Id A law article has a unique law article id corresponding to it
Agree With A case agree with several pieces of key information
Applicable Law The law mentioned in the judgment documents that matches the case
Occur in Time The verdict in a case occurs on a fixed date
Reason Case Documents contain the reason of lawsuit in the case
Detail The name of a case corresponds to the specific information of the case
As shown in Figure 2(a), the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG). The LAKG includes original article of law, law article id, and key information. The key information is extracted from the original article of law using a LLM. For a more detailed instruction of LLM prompts in the task of retrieving law article key information or a specific example, we direct the reader to Section 1.1 of supplementary material. The ACKG contains the name of a case, time of the court session, prosecution reason, and the specifics of a case. The description of the case process is derived from the content of the judgment documents processed using a LLM. For a more detailed instruction of LLM prompts in the task of retrieving case process, we direct the reader to Section 1.2 of supplementary material.
Based on the reference law articles written in the judgment documents, it is possible to link the name of a case node with law article id nodes. We analyze the correlation between case information and the corresponding key information of applicable law articles using a LLM and ultimately select up to five of the most relevant keywords for connection. For a more detailed instruction of LLM prompts in the task of matching case information to the corresponding key information of applicable law articles, we direct the reader to Section 1.3.2 of supplementary material. To ensure the consistency and accuracy of the knowledge graph. We ignore any keywords that do not exist in the graph output by the large language model. Tables 1 and 2 provide the details of all types of nodes and relationships in the CLAKG.
2.2 LLM and CLAKG Based Law Article Recommendation Framework
Figure 3: The LLM and CLAKG based law article recommendation framework.
The LLM and CLAKG based law article recommendation framework is shown in figure 3. Before the start of the law article recommendation task, the CLAKG constructed in Section 2.1 needs to undergo graph embedding preprocessing using the Relational Graph Convolutional Network (RGCN) model mentioned in Section 2.2.1. The process of law article recommendation begins with the user inputting a newly emerged case. In Section 2.2.2, a LLM is used to match the k most relevant keywords to this new case (with k=8 as selected in this paper). Based on this, in Section 2.2.3, a candidate law article retrieval algorithm is employed to retrieve q candidate law articles that may be suitable for the new case (with q=5 as chosen in this paper). In Section 2.2.4, the information from the new case input by the user is integrated with the candidate law articles and historical case information obtained in Section 2.2.3 into a prompt, which is then passed to the LLM to complete the law article recommendation task.
2.2.1 Graph Embedding on CLAKG with RGCN
Graph embedding involves utilizing the topological structure information in knowledge graphs to assign embedding vectors to each node within the graph. The Graph Convolutional Network(GCN) model, initially proposed by [47], was the first graph neural network used for graph embedding. [55] improved upon the GCN, introducing the Relation Graph Convolution Neural Networks (RGCN), which assigns different weights to different relationships, enhancing the quality of graph embeddings.
The node embedding vector update formula for the RGCN model is shown in Equation 1.
$h_{i}^{(l+1)}=\sigma\left(\sum_{r\in\mathcal{R}}\sum_{j\in\mathcal{N}_{i}^{r}}\frac{1}{c_{i,r}}W_{r}^{(l)}h_{j}^{(l)}+W_{0}^{(l)}h_{i}^{(l)}\right)$ (1)
where $h_{i}^{(l)}$ represents the embedding vector of node $i$ in the $l$-th layer, $N_{i}^{r}$ represents all neighboring nodes of node $i$ that have relationship $r$. $c_{i,r}$ is a hyperparameter, which in this paper is taken as $\left|N_{i}^{r}\right|$. Both $W_{r}^{(l)}$ and $W_{0}^{(l)}$ are parameter matrices.
The graph embedding training task adopted in this paper is link prediction. For the link prediction task, part of the dataset consists of several triples (head node, relation, tail node) selected from the knowledge graph $G=(\mathcal{V},\mathcal{E},\mathcal{R})$, which are used as positive examples. Another part of the dataset consists of randomly selected entities and relations from the knowledge graph, which are used to construct triples (head node, relation, tail node) that do not actually have a connection in the graph, i.e., fabricated triples, used as negative examples. The training objective of the model is to distinguish whether the given triples truly exist in the graph or are fabricated. The number of positive and negative triples is equal, which helps to prevent bias during model training.
Algorithm 1 Candidate Law Article Retrieval Algorithm
1:key information node, the embedded vector dictionary obtained in Section 2.2.1
2:5 Original articles of law node
1:Initialize: law_article_distances = [] , law_article_set = {} # Initializes a list to store distances and corresponding law articles and a set to store candidate law articles
2:for $e\in\text{key information node}$ do
3:  Case_set.add_node($x\mid x-[\text{key}]\rightarrow e$) # $x$ is ‘Original article of penal law’ node
4:end for
5:for $\text{law\_article}\in\text{law\_article\_set}$ do
6:  $\text{distance}=0$ # Initializes the distance score
7:  for $e\in\text{key information node}$ do
8:   $\text{distance}\mathrel{+}=\operatorname{dis}(\text{emb}[\text{law\_article}],\text{emb}[\text{e}])$ # Calculate the distance score
9:    where $\operatorname{dis}(a,b)=\frac{a\cdot b}{\|a\|\times\|b\|}$ # Cosine similarity formula
10:  end for
[…]
12:end for
[…]
14:return law_article_distances[:5] # Output the top five candidate law articles
Figure 4: Flowchart of RGCN training process.
The training process of the model is shown in Figure 4. Firstly, the system randomly initializes the embedding vectors for all nodes and relations. Based on this, the system updates the node embeddings using Equation 1. The system then calculates the score based on the DisMult decomposition, as described in Equation 2. This paper selects the cross-entropy function as the loss function, as described in Equation 3. When the triple is a positive example, meaning the information truly exists in the knowledge graph, the value of $y$ is 1. When the triple is a negative example, meaning the information does not exist in the graph, the value of $y$ is 0 .
$f(s,r,o)=e_{s}^{T}R_{r}e_{o}$ (2)
where $e_{s}$ and $e_{o}$ represent the embedding vectors of the nodes s and o, and $R_{r}$ is the embedding vector of the relation r.
$\operatorname{Loss}=$ $-\frac{1}{(1+\omega)|\hat{\varepsilon}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Bigg[y\log l(f(s,r,o))$ (3)
$+(1-y)\log\left(1-l(f(s,r,o))\right)\Bigg]$
where $\mathcal{J}$ represents all the datasets, and $l$ is the logistic activation function, as described in Equation 4.
$l(x)=\frac{1}{1+e^{-x}}=\frac{e^{x}}{e^{x}+1}$ (4)
2.2.2 Matching Key Information Nodes in CLAKG with LLM
In this section, we aim to match the relevant key information nodes within CLAKG based on current case information provided by users. To facilitate this process, we have designed prompts for the LLM, which include the expert role, task description, details of the newly reported case, known key information nodes, and example outputs. For a more detailed instruction of LLM prompts in the task of matching key information nodes in the CLAKG related to new case information, we direct the reader to Section 2.1 of supplementary material.
2.2.3 Law Article and Case Retrieval Algorithm
In this section, we propose an algorithm for retrieving historical case information and candidate law articles by leveraging the aforementioned data to identify the law article most relevant to the current case.
The proposed method for retrieving candidate law article involves traversing all ‘key information’ nodes obtained in 2.2.2 and retrieving the ‘Original article of law’ nodes that are connected to the ‘key information’ nodes via ‘key’ relationships. These nodes are then aggregated into a set container, with all ‘Original article of law’ nodes in the set representing the candidate law article. Given that the number of candidate law articles in the set container may be substantial, it is essential to select those law articles most analogous to the newly reported case for further analysis. To this end, each node in the container is evaluated by calculating the distance scores (cosine similarity) between the ‘Original article of penal law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 2.2.1.
These distance scores are summed to yield a cumulative distance score between each ‘Original article of law’ node and the newly reported case. Ultimately, the top five ‘Original article of law’ nodes with the highest cumulative distance scores are selected as the candidate law articles, serving as reference points for the LLM in completing the law article recommendation task.
The pseudo-code is shown in Algorithm 1. The time complexity of this algorithm is $O(m\cdot n)+O(m\log m)$, where $m$ is the number of elements in the law article set and $n$ is the number of elements in the key information node set.
After the above process, by querying the CLAKG for nodes representing ‘the name of a case’ connected to these law articles, we can obtain historical case information for reference.
2.2.4 LLM-Enhanced Law Article Recommendation
Our method intends to incorporate reference case information and the original text of candidate law articles to eliminate the hallucinations of LLMs. After retrieving candidate law articles and relevant historical cases with the method described in Section 2.2.3, The knowledge graph is then queried for key details, and this integrated information helps determine the applicability of the candidate articles. The prompt used for the LLM in this chapter consists of expert roles, task descriptions, information on new criminal cases, reference case information, candidate law articles, and output examples.
For a more detailed instruction of LLM prompts in the task of law article recommendation, we direct the reader to Section 2.3 of supplementary material. After the user describes the case information to the system, the system returns the results generated by the LLM. The user can further inquire about the results provided by the LLM.
3 Experiment
3.1 Datasets
Figure 5: The name of a case, time of the court session, prosecution reason on the China Judgments Online.
The method proposed in this paper can be applied to any type of law articles in any country. For research convenience, the law articles datasets employed in this section is derived from the Amendment (XI) to the Criminal Law of the People’s Republic of China [57], ratified during the 24th session of the Standing Committee of the 13th National People’s Congress. This amendment is systematically structured into two sections: the General Provisions and the Specific Provisions, encompassing a total of 452 articles—101 in the General law articles and 351 in the Specific law articles. The CLAKG was constructed utilizing the methodology in Section 2.1.
The criminal case data analyzed in this study were obtained from the China Judgments Online [58] platform. Figure 5 presents one of the cases. In light of the fact that the Amendment (XI) to the Criminal Law of the People’s Republic of China came into force on March 1, 2021, all judicial decisions used herein pertain to cases adjudicated after this date. During this period, due to concerns over information security and privacy protection, the number of publicly available criminal case judgments was limited. As a result, certain law articles corresponded with only a limited number of cases, while some provisions had no corresponding cases at all. We curated a dataset comprising 1,170 case records, which collectively address law articles from 85 articles within the Specific Provisions. To optimize the dataset, we divided it in a manner that maximized the diversity of law articles represented in the test set. Specifically, 997 case records were designated for the training set and 173 for the test set. The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zoro_shot_num = 36). The Adjudicated Case Knowledge Graph was subsequently constructed based on the training data, following the approach outlined in Section 2.1.
3.2 CLAKG Construction and Graph Embedding
We employed the method described in Section 2.1 to integrate the LAKG and the ACKG, resulting in the CLAKG. The types and numbers of nodes in CLAKG are presented in Table 3, while the types and numbers of relationships are detailed 4.
Table 3: Entity types and numbers in CLAKG.
Entity Type Number
Original article of law 452
Key information 3405
Law article id 452
The name of a case 997
Time of the court session 997
Prosecution reason 997
The specifics of the case 997
Table 4: Relation types and numbers in CLAKG.
Relation Type Number
Key 7854
Id 452
Agree with 997
Applicable law 4118
Occur in time 997
Reason 997
Detail 997
We employed the RGCN model described in Section 2.2.1 to perform graph embedding preprocessing on CLAKG. Our hyperparameter choices were as follows: $h_{dim}=16$, test_size=0.2test\_size=0.2, learning_rate=0.01learning\_rate=0.01, num_epochs=50num\_epochs=50. We selected the model results corresponding to the highest Test AUC as the final graph embedding outcome. The Train AUC and Test AUC curves are presented in Figure 6. Upon examining this figure, we observed that the Train AUC exhibited an overall upward trend. The Test AUC increased when $Epoch\leq 31$ and decreased when $Epoch\geq 31$. Consequently, we selected the model results at Epoch 31 as the final graph embedding outcome.
Figure 6: Train loss and AUC over epochs.
3.3 Case Study
We illustrate the law recommendation process based on LLM and CLAKG using the case of ”Zhang Yue’s offenses of bribery and abuse of power” as an example. Initially, the user inputs the details of a new case:
During Zhang Yue’s tenure as cultural administrator and grid member of Yongfeng Village, Zhang Yue manipulated residential information to secure illicit benefits, accepting 72,500 yuan in bribes. After the crime was uncovered, he voluntarily admitted his wrongdoing and returned the funds.
The system then retrieves keywords pertinent to the case using the methodology outlined in Section 2.2.2, identifying the following terms: [‘accepting bribes’, ‘abuse of power’, ‘bribery’]. Subsequently, the system locates the corresponding node IDs within the CLAKG and obtains the graph embedding vectors for these nodes as described in Section 3.2. Utilizing the historical case information and the candidate law articles retrieval algorithm specified in Section 2.2.3, the system identifies the candidate law articles and their associated cases, as depicted in Section 2.2 of supplementary material.
The system consolidates the above information into a prompt using the method described in Section 2.2.4, and submits it to the LLM to carry out the law recommendation task. Ultimately, the LLM outputs “Article 385” as the recommended law article. The result produced by the LLM is consistent with the outcome recorded in the court judgment.
3.4 Test Result
3.4.1 Analyzing the Impact of Label Imbalance in Training
We utilized the BERT model for law article recommendation as comparison methods. This method was compared against the method proposed in this paper: the LLM(specifically OpenAI’s ChatGPT-4.0) for law article recommendation based on CLAKG (with case information).
Table 5: Model Accuracy Comparison Between Bert and Ours
Model Accuracy
BERT (baseline) 0.289
LLM+CLAKG (OURS, with case) 0.694
As shown in the Table 5, the accuracy of the BERT model is only 0.289. Upon analyzing the results provided by the BERT method, we discovered that this method classified all test cases under ”Article 133.” We believe this could be due to the small size of the training set, an imbalance in training labels [42], or an inappropriate loss function. Even after replacing the cross-entropy loss function with the Focal Loss function, the model still classified all test cases as ”Article 133”.
The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than that of the BERT-based method. This indicates that the proposed approach effectively mitigates the impact of the insufficient data and label imbalance in training.
3.4.2 Analyzing the Impact of Hallucinations in LLMs
We utilized the LLM (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods) for law article recommendation as comparison method. This method was compared against the method proposed in this paper: the LLM for law article recommendation based on CLAKG (with case information).
Table 6: Model Accuracy Comparison Between LLM and Ours
Model Accuracy
LLM (baseline) 0.549
LLM+CLAKG (OURS, with case) 0.694
As shown in the Table 6, the accuracy of the LLM without external legal data is significantly higher than that of the BERT model, achieving an accuracy of 0.549. However, it is still lower than the proposed model (LLM+CLAKG). We attribute this to the hallucinations of LLM and the proposed approach effectively mitigates the impact of label imbalance in training. The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than that of the LLM-based method. This indicates that the proposed approach effectively mitigates the Huallucinations in LLMs.
3.4.3 Analyzing the Impact of Utilizing Macro Semantics
We utilized the LLM (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods) based on TFIDF for law article recommendation as comparison method. This method was compared against two method proposed in this paper: the LLM for law article recommendation based on CLAKG (with case information) and the LLM for law article recommendation based on CLAKG (no case information).
Table 7: Model Accuracy Comparison Between LLM+TFIDF and Ours
Model Accuracy
LLM+TFIDF (base line) 0.595
LLM+CLAKG (OURS, no case) 0.676
LLM+CLAKG (OURS, with case) 0.694
As shown in the Table 7, both the TFIDF-based law article recommendation method and the CLAKG-based law article recommendation method mitigate the hallucinations of the LLM and improve its accuracy, with the latter performing better. We believe this is because the former only matches at the text level, while the latter matches the new case with keywords at the semantic level. Additionally, the latter utilizes all information within the CLAKG through graph embedding techniques, allowing it to calculate the most relevant law articles from a holistic perspective.
4 Conclusion
This paper proposes an efficient law article recommendation approach using a Case-Enhanced Law Article Knowledge Graph (CLAKG) combined with a Large Language Model (LLM). CLAKG integrates law articles and case information. It is characterized by a rich topological structure formed through usage relationships and shared key information found in judgments, which enhances the effectiveness of law article recommendation tasks. The proposed approach integrates LLMs with CLAKG, enabling more accurate recommendations of law articles and related cases by utilizing macro semantics to mitigate LLM hallucinations. The effectiveness of the proposed approach is verified through comprehensive comparisons with several baseline models on the law article recommendation task. This approach effectively addresses challenges such as insufficient data, imbalanced labels, and LLM hallucinations, leading to a significant improvement in the accuracy of the law article recommendation task.
In future work, we plan to expand the data volume of CLAKG, particularly by incorporating more case data to further enhance the performance of law article recommendation. Additionally, we will explore further applications of the CLAKG-LLM integration to improve the efficiency of court judgments.

### v2

@@
 Finally, through a series of experiments using judgment documents from the website ”China Judgements Online”, we have improved the accuracy of law article recommendation in cases from 0.549 to 0.694, demonstrating that our proposed method significantly outperforms baseline approaches.
+keywords:
+law article recommendation , Large Language Model , Knowledge Graph
+[…]
+[label1]organization=Global Institute of Future Technology, Shanghai Jiao Tong University, city=Shanghai, postcode=200240, state=Shanghai, country=China
+[…]
+[label2]organization=Donghua University, city=Shanghai, postcode=201620, state=Shanghai, country=China
+{highlights}
+Proposed a novel Case-Enhanced Law Article Knowledge Graph (CLAKG) that integrates law articles and historical case data, enhancing the accuracy of law article recommendations.
+Introduced an automated CLAKG construction method that uses a Large Language Model (LLM) to reduce manual input and improve scalability.
+Developed a closed-loop recommendation system, leveraging LLMs and CLAKG for more accurate law article recommendations while mitigating common issues like hallucinations in LLM outputs.
+Achieved significant improvement in accuracy: The proposed method improved law article recommendation accuracy from 0.549 to 0.694, outperforming baseline models such as BERT, DPCNN, TFIDF-RAG, Graph-RAG and Light-RAG.
 1 Introduction
-Figure 1: The pipeline of CLAKG. The Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG).
-In a modern rule-of-law society, the efficiency of court judgments is crucial for ensuring economy growth and social stability [39, 53, 41].
+In a modern rule-of-law society, the efficiency of court judgments is crucial for ensuring economic growth and social stability[1, 2, 3].
 Meanwhile, [4] revealed that the current backlog of cases in courts is enormous, which means the workload of judicial personnel is heavy.
@@
 Therefore, there is a need to solve above problems.
-Some researchers utilized Large Language Model (LLM) to complete the task of law article recommendation [61, 50, 56].
-When engaging with extensive LLMs, users are enabled to not only retrieve the numeral and textual format of law article but also to investigate the rationale behind the applicability of these statutes to the specific case by posing supplemental inquiries.
+Some researchers utilized Large Language Models (LLMs) to complete the task of law article recommendation[12, 13, 14].
+By using a large LLM, users can not only retrieve the law article in numeric and textual form, but also investigate the rationale behind its applicability to the case by asking follow-up questions.
 Moreover, users have the capability to probe into the rationale for the non-selection of other analogous provisions.
@@
 To demonstrate the effectiveness of our proposed method, we constructed a dataset of Chinese criminal law judgments.
-We compared the method introduced in this paper with approaches based on BERT, methods utilizing LLM, and the LLM approach based on TFIDF.
+We compared the method introduced in this paper with approaches based on BERT[23], GRU[8], DPCNN[11], and several RAG approaches such as TFIDF-RAG[21], Graph-RAG[27] and Light-RAG[28].
+Finally, we conducted an ablation study of our method.
 The remainder of this paper is organized as follows.
+Section 2 reviews related works about law article recommendation, RAG technology and knowledge graphs in legal domain.
 Section 3 proposes a closed-loop law article recommendation framework.
-Section 3 utilizes the method from the second section to construct the CLAKG and compares the accuracy of this method with other baseline approaches.
+Section 4 utilizes the method from the third section to construct the CLAKG and compares the accuracy of this method with other baseline approaches.
 Lastly, the conclusions and future work are highlighted in Section 5.
+2 Related Work
+2.1 Law Article Recommendation
+Early approaches to law article recommendation treated the task as a text classification or retrieval problem.
+Classical methods relied on keyword matching or manual feature engineering, but these struggled with the complex language of statutes.
+With the rise of deep learning, researchers applied neural models like LSTMs[6] and CNNs[7] to match case fact descriptions with relevant law articles.
+These models improved over bag-of-words baselines by capturing sequential and local context features.
+Transformer-based language models further advanced the field of BERT[29] and its legal-domain adaptations (e.g., LegalBERT[30]) enabled richer semantic representations of legal text.
+For example, Chalkidis et al.[31] showed that a BERT-based classifier can predict applicable European human rights articles more accurately than traditional models, while Aletras et al.[32] demonstrated early on the feasibility of data-driven prediction of court decisions and their pertinent articles.
+Despite these gains, most prior models formulate law article recommendation as mapping facts to article IDs, without directly leveraging the content of the law articles themselves.
+This focus solely on fact-to-ID correspondence makes them heavily data-dependent and sensitive to label imbalance in training sets[33].
+In practice, it also limits interpretability – the models do not explain why a recommended article is relevant, since the law articles are not explicitly used in the reasoning process.
+These limitations have motivated explorations into techniques that can incorporate external legal knowledge and provide better justification for recommendations.
+2.2 LLMs and Retrieval-Augmented Generation Techniques
+The advent of large language models (LLMs), such as GPT-style transformers, has opened new avenues for law article recommendation.
+Rather than training a specific classifier, one can prompt an LLM with a case’s facts and ask for relevant law articles.
+Recent works have started to investigate this approach[12, 13, 14], noting that LLMs can, in principle, leverage their vast pretrained knowledge to identify relevant law articles and even generate explanations.
+In particular, Shui et al.[14] evaluate various state-of-the-art LLMs on legal judgment prediction subtasks (including relevant law article prediction), finding that while LLMs exhibit promising zero-shot capabilities, they still lag behind task-specific models on accuracy.
+A major concern with directly using LLMs for legal guidance is hallucination: the model may produce plausible-sounding but incorrect or nonexistent legal references.
+Dahl et al.[15] provide a sobering analysis of this issue, showing that even advanced LLMs frequently fabricate legal citations or statutes, which is unacceptable in high-stakes domains like law.
+To mitigate such hallucinations, researchers are turning to Retrieval-Augmented Generation (RAG) techniques.
+RAG combines LLMs with information retrieval: before or during text generation, the system retrieves relevant documents (e.g.
+law articles or case precedents) and conditions the LLM on that retrieved evidence.
+By grounding responses in actual law articles, RAG can significantly improve factual accuracy and trustworthiness of the model’s output.
+This approach was originally popularized in open-domain QA systems[16] and has since been applied to legal NLP as well.
+For instance, a RAG-based legal assistant might fetch the text of candidate law articles given a case description, and the LLM then quotes or summarizes those law articles when recommending applicable law, thereby rooting its answer in verifiable sources.
+Key components of RAG include a retriever and a generator.
+In the legal domain, classic IR methods like TF–IDF[21] or BM25[22] have been used to retrieve statute candidates by keyword overlap, while more recent approaches use neural retrievers (e.g.
+dense embeddings).
+Borgeaud et al.[23] showed that even a frozen BERT-based retriever can substantially improve generation quality by supplying relevant text from a large corpus.
+However, naive retrieval based on surface text may miss conceptual matches.
+Legal phrasing varies, and a model might need to retrieve statutes using synonymous terms or implied connections, something neural retrieval aims to handle better than lexical methods.
+Overall, RAG represent a promising direction for law article recommendation, as they combine the reasoning ability of generative models with the grounding of information retrieval, thereby reducing hallucinations and improving the transparency of the recommendation.
+2.3 Knowledge Graphs in the Legal Domain
+Beyond unstructured text retrieval, Knowledge Graphs (KGs) have emerged as valuable resources for legal information retrieval and recommendation.
+A legal knowledge graph represents entities such as law articles, cases, legal concepts, and their relationships (e.g., a case cites a statute, or a chapter includes certain articles) in a structured form.
+This graph structure can complement text-based methods by capturing the domain’s ontology and the interconnections between law articles.
+Prior works have constructed legal knowledge graphs to assist various tasks.
+For example, law-specific knowledge bases have been used to integrate charge definitions and article relationships into prediction models, yielding performance gains and better interpretability in tasks like judgment prediction[34, 35].
+More recently, Chaudhri et al.[24] argued that KGs are well-suited as the underlying database for RAG systems: instead of retrieving free text, one can query a knowledge graph to retrieve facts or linked nodes, ensuring that the generation is grounded in a consistent knowledge structure.
+In the legal domain, a KG could be queried to find statutes related to certain legal concepts or past cases with similar fact patterns.
+One advantage of KGs is that they store high-level associations beyond lexical similarity – for instance, a graph can link a section of law to broader categories or to precedent cases, enabling the recommendation system to recall relevant law articles even if a case’s description uses different wording.
+There is growing interest in integrating KGs with LLMs[25].
+Pan et al.
+[25] outline a roadmap for unifying the strengths of symbolic knowledge graphs and distributed language models, noting that KGs can provide factual grounding for LLMs while LLMs can help populate and expand KGs from raw texts.
+In the context of law, researchers have begun designing hybrid systems that use KGs alongside LLMs.
+Chen et al.[33], for instance, construct a case-enhanced law article KG that links statutes with relevant past cases, and then use an LLM-based pipeline to recommend articles by traversing the graph for pertinent nodes (law articles) and using the LLM to generate the final recommendations and explanations.
+Such approaches have demonstrated improved accuracy over text-only baselines, highlighting that the structured knowledge in KGs can effectively guide the retrieval and generation process.
+Meanwhile, Edge et al.[26] explore automatic construction of knowledge graphs using LLMs (generating graph triples from unstructured legal text) and using those graphs to enhance answer generation.
+This suggests a future where legal KGs and LLMs continuously enrich each other: the KG offers a source of truth to reduce LLM errors, and the LLM helps keep the KG up-to-date with new insights.
+The integration of knowledge graphs in legal AI is still an emerging area, and challenges remain, such as ensuring the completeness and currency of the graph.
+Nonetheless, the combination of KGs with retrieval-augmented LLMs holds great promise for building more reliable and interpretable law article recommendation systems in the coming years.
 3 Method
+Figure 1: The pipeline of CLAKG. The Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG).
 The pipeline of CLAKG-based law article recommendation is shown in Figure 1.
@@
 3.1 Case-Enhanced Law Article Knowledge Graph Construction
-(a) Schema of CLAKG. Ellipses of different colors represent
+Figure 2: Schema of CLAKG. Ellipses of different colors represent
 different entity types.
-(b) Example of CLAKG.
-Figure 2: Overall schema and instantiation example.
+Figure 3: Example of CLAKG.
 Table 1: Entity types in Case-Enhanced Law Article Knowledge Graph
@@
 Detail The name of a case corresponds to the specific information of the case
-As shown in Figure 2(a), the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG).
+As shown in Figure 2, the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG).
 The LAKG includes original article of law, law article id, and key information.
 The key information is extracted from the original article of law using a LLM.
-For a more detailed instruction of LLM prompts in the task of retrieving law article key information or a specific example, we direct the reader to Section 1.1 of supplementary material.
 The ACKG contains the name of a case, time of the court session, prosecution reason, and the specifics of a case.
 The description of the case process is derived from the content of the judgment documents processed using a LLM.
-For a more detailed instruction of LLM prompts in the task of retrieving case process, we direct the reader to Section 1.2 of supplementary material.
 Based on the reference law articles written in the judgment documents, it is possible to link the name of a case node with law article id nodes.
 We analyze the correlation between case information and the corresponding key information of applicable law articles using a LLM and ultimately select up to five of the most relevant keywords for connection.
-For a more detailed instruction of LLM prompts in the task of matching case information to the corresponding key information of applicable law articles, we direct the reader to Section 1.3.2 of supplementary material.
-To ensure the consistency and accuracy of the knowledge graph.
-We ignore any keywords that do not exist in the graph output by the large language model.
+To ensure the consistency and accuracy of the knowledge graph, we ignore any keywords that do not exist in the graph output by the large language model.
 Tables 1 and 2 provide the details of all types of nodes and relationships in the CLAKG.
@@
 The number of positive and negative triples is equal, which helps to prevent bias during model training.
-Algorithm 1 Candidate Law Article Retrieval Algorithm
-1:key information node, the embedded vector dictionary obtained in Section 2.2.1
-2:5 Original articles of law node
-1:Initialize: law_article_distances = [] , law_article_set = {} # Initializes a list to store distances and corresponding law articles and a set to store candidate law articles
-2:for $e\in\text{key information node}$ do
-3:  Case_set.add_node($x\mid x-[\text{key}]\rightarrow e$) # $x$ is ‘Original article of penal law’ node
-4:end for
-5:for $\text{law\_article}\in\text{law\_article\_set}$ do
-6:  $\text{distance}=0$ # Initializes the distance score
-7:  for $e\in\text{key information node}$ do
-8:   $\text{distance}\mathrel{+}=\operatorname{dis}(\text{emb}[\text{law\_article}],\text{emb}[\text{e}])$ # Calculate the distance score
-9:    where $\operatorname{dis}(a,b)=\frac{a\cdot b}{\|a\|\times\|b\|}$ # Cosine similarity formula
-10:  end for
-[…]
-12:end for
-[…]
-14:return law_article_distances[:5] # Output the top five candidate law articles
 Figure 5: Flowchart of RGCN training process.
@@
 where $e_{s}$ and $e_{o}$ represent the embedding vectors of the nodes s and o, and $R_{r}$ is the embedding vector of the relation r.
-$\operatorname{Loss}=$ $-\frac{1}{(1+\omega)|\hat{\varepsilon}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Bigg[y\log l(f(s,r,o))$ (3)
-$+(1-y)\log\left(1-l(f(s,r,o))\right)\Bigg]$
+$\operatorname{Loss}=$ $-\frac{1}{(1+\omega)|\hat{\varepsilon}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Bigg[y\log l(f(s,r,o))$ $+(1-y)\log\left(1-l(f(s,r,o))\right)\Bigg]$ (3)
 where $\mathcal{J}$ represents all the datasets, and $l$ is the logistic activation function, as described in Equation 4.
@@
 To facilitate this process, we have designed prompts for the LLM, which include the expert role, task description, details of the newly reported case, known key information nodes, and example outputs.
-For a more detailed instruction of LLM prompts in the task of matching key information nodes in the CLAKG related to new case information, we direct the reader to Section 2.1 of supplementary material.
+The prompts for LLM in the task of matching key information nodes in the CLAKG related to new case information is shown in Table 3.
+Table 3: Prompts for LLM in the task of matching key information nodes in the CLAKG related to new case information.
+Component Details
+Expert Role Expert in law article analysis
+Task Description Given the available new case information and the following key information nodes, please output 080-8 key information nodes most relevant to the case.
+If you believe there are fewer than 8 relevant key information nodes, there is no need to force the list to reach 8 nodes.
+New Case Information Description of the new case input by the user
+Key Information Nodes all key information nodes in the CLAKG
+Precautions Please directly output the key information nodes, separated by semicolons (;) without including any additional content (including unnecessary punctuation marks).
+Output Example Public and private property; multiple thefts; sale; committing a crime; causing damage; seriously disrupting public order; multiple thefts; lawfully performing duties
 3.2.3 Law Article and Case Retrieval Algorithm
+Algorithm 1 Candidate Law Article Retrieval Algorithm
+1:key information node, the embedded vector dictionary obtained in Section 3.2.1
+2:5 Original articles of law node
+1:Initialize: law_article_distances = [] , law_article_set = {} $\triangleright$ Initializes a list to store distances and corresponding law articles and a set to store candidate law articles
+2:for $e\in\text{key information node}$ do
+3:  Case_set.add_node($x\mid x-[\text{key}]\rightarrow e$) $\triangleright$ $x$ is ‘Original article of penal law’ node
+4:end for
+5:for $\text{law\_article}\in\text{law\_article\_set}$ do
+6:  $\text{distance}=0$ $\triangleright$ Initializes the distance score
+7:  for $e\in\text{key information node}$ do
+8:   $\text{distance}\mathrel{+}=\operatorname{dis}(\text{emb}[\text{law\_article}],\text{emb}[\text{e}])$ $\triangleright$ Calculate the distance score
+9:    where $\operatorname{dis}(a,b)=\frac{a\cdot b}{\|a\|\times\|b\|}$ $\triangleright$ Cosine similarity formula
+10:  end for
+[…]
+12:end for
+[…]
+14:return law_article_distances[:5] $\triangleright$ Output the top five candidate law articles
 In this section, we propose an algorithm for retrieving historical case information and candidate law articles by leveraging the aforementioned data to identify the law article most relevant to the current case.
-The proposed method for retrieving candidate law article involves traversing all ‘key information’ nodes obtained in 2.2.2 and retrieving the ‘Original article of law’ nodes that are connected to the ‘key information’ nodes via ‘key’ relationships.
+The proposed method for retrieving candidate law article involves traversing all ‘key information’ nodes obtained in 3.2.2 and retrieving the ‘Original article of law’ nodes that are connected to the ‘key information’ nodes via ‘key’ relationships.
 These nodes are then aggregated into a set container, with all ‘Original article of law’ nodes in the set representing the candidate law article.
 Given that the number of candidate law articles in the set container may be substantial, it is essential to select those law articles most analogous to the newly reported case for further analysis.
-To this end, each node in the container is evaluated by calculating the distance scores (cosine similarity) between the ‘Original article of penal law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 2.2.1.
+To this end, each node in the container is evaluated by calculating the distance scores (cosine similarity) between the ‘Original article of penal law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 3.2.1.
 These distance scores are summed to yield a cumulative distance score between each ‘Original article of law’ node and the newly reported case.
@@
 The prompt used for the LLM in this chapter consists of expert roles, task descriptions, information on new criminal cases, reference case information, candidate law articles, and output examples.
-For a more detailed instruction of LLM prompts in the task of law article recommendation, we direct the reader to Section 2.3 of supplementary material.
-After the user describes the case information to the system, the system returns the results generated by the LLM.
-The user can further inquire about the results provided by the LLM.
+After the user describes the case information to the system, the system returns the results generated by the LLM. The user can further inquire about the results provided by the LLM.
 4 Experiment
@@
 Specifically, 997 case records were designated for the training set and 173 for the test set.
-The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zoro_shot_num = 36).
+The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zoro shot number = 36).
 The Adjudicated Case Knowledge Graph was subsequently constructed based on the training data, following the approach outlined in Section 3.1.
 4.2 CLAKG Construction and Graph Embedding
-We employed the method described in Section 2.1 to integrate the LAKG and the ACKG, resulting in the CLAKG.
-The types and numbers of nodes in CLAKG are presented in Table 3, while the types and numbers of relationships are detailed 4.
-Table 3: Entity types and numbers in CLAKG.
-Entity Type Number
-Original article of law 452
-Key information 3405
-Law article id 452
-The name of a case 997
-Time of the court session 997
-Prosecution reason 997
-The specifics of the case 997
-Table 4: Relation types and numbers in CLAKG.
-Relation Type Number
-Key 7854
-Id 452
-Agree with 997
-Applicable law 4118
-Occur in time 997
-Reason 997
-Detail 997
+We employed the method described in Section 3.1 to integrate the LAKG and the ACKG, resulting in the CLAKG. The types and numbers of nodes as well as relationships in CLAKG are presented in Table 4.
+Table 4: Entity and relation types with their Numbers in CLAKG.
+Entity Type Number Relation Type Number
+Original article of law 452 Key 7854
+Key information 3405 Id 452
+Law article id 452 Agree with 997
+The name of a case 997 Applicable law 4118
+Time of the court session 997 Occur in time 997
+Prosecution reason 997 Reason 997
+The specifics of the case 997 Detail 997
+Figure 7: Train loss and AUC over epochs.
 We employed the RGCN model described in Section 3.2.1 to perform graph embedding preprocessing on CLAKG.
@@
 Consequently, we selected the model results at Epoch 31 as the final graph embedding outcome.
-Figure 6: Train loss and AUC over epochs.
 4.3 Case Study
-We illustrate the law recommendation process based on LLM and CLAKG using the case of ”Zhang Yue’s offenses of bribery and abuse of power” as an example.
+We illustrate the law article recommendation process based on LLM and CLAKG using the case of ”Zhang Yue’s offenses of bribery and abuse of power” as an example.
 Initially, the user inputs the details of a new case:
@@
 Subsequently, the system locates the corresponding node IDs within the CLAKG and obtains the graph embedding vectors for these nodes as described in Section 4.2.
-Utilizing the historical case information and the candidate law articles retrieval algorithm specified in Section 2.2.3, the system identifies the candidate law articles and their associated cases, as depicted in Section 2.2 of supplementary material.
-The system consolidates the above information into a prompt using the method described in Section 2.2.4, and submits it to the LLM to carry out the law recommendation task.
+Utilizing the historical case information and the candidate law articles retrieval algorithm specified in Section 3.2.3, the system identifies the candidate law articles and their associated cases.
+The system consolidates the above information into a prompt using the method described in Section 3.2.4, and submits it to the LLM to carry out the law article recommendation task.
 Ultimately, the LLM outputs “Article 385” as the recommended law article.
@@
 4.4.1 Analyzing the Impact of Label Imbalance in Training
-We utilized the BERT model for law article recommendation as comparison methods.
-This method was compared against the method proposed in this paper: the LLM(specifically OpenAI’s ChatGPT-4.0) for law article recommendation based on CLAKG (with case information).
-Table 5: Model Accuracy Comparison Between Bert and Ours
+We utilized the BERT[23], GRU[8] and DPCNN[11] model for law article recommendation as comparison methods.
+These methods were compared against the method proposed in this paper: the LLM(specifically OpenAI’s ChatGPT-4.0) for law article recommendation based on CLAKG (with case information).
+Table 5: Model accuracy comparison between classical models and ours
 Model Accuracy
 BERT (baseline) 0.289
-LLM+CLAKG (OURS, with case) 0.694
-As shown in the Table 5, the accuracy of the BERT model is only 0.289.
+GRU (baseline) 0.427
+DPCNN (baseline) 0.495
+LLM + CLAKG (OURS, no case) 0.676
+As shown in the Table 5, the accuracy of the BERT (baseline) is only 0.289.
 Upon analyzing the results provided by the BERT method, we discovered that this method classified all test cases under ”Article 133.” We believe this could be due to the small size of the training set, an imbalance in training labels[40], or an inappropriate loss function.
 Even after replacing the cross-entropy loss function with the Focal Loss function, the model still classified all test cases as ”Article 133”.
-The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than that of the BERT-based method.
+The GRU (baseline) model achieved an accuracy of 0.427.
+GRU, or Gated Recurrent Unit, is a type of recurrent neural network (RNN) that is widely used for sequential data, and this result represents its performance on the task at hand.
+The DPCNN (baseline) model, which stands for Deep Pyramid Convolutional Neural Network, performed slightly better, with an accuracy of 0.495.
+DPCNN is a convolutional neural network architecture designed for text classification tasks, and its accuracy reflects its ability to capture hierarchical patterns in the input data.
+The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than any one of the classic methods.
 This indicates that the proposed approach effectively mitigates the impact of the insufficient data and label imbalance in training.
 4.4.2 Analyzing the Impact of Hallucinations in LLMs
-We utilized the LLM (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods) for law article recommendation as comparison method.
-This method was compared against the method proposed in this paper: the LLM for law article recommendation based on CLAKG (with case information).
-Table 6: Model Accuracy Comparison Between LLM and Ours
+We utilized the LLM[41] (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods), TFIDF-RAG[21], Graph-RAG[27] and Light-RAG[28] for law article recommendation as comparison method.
+This method was compared against the method proposed in this paper: the LLM for law article recommendation based on CLAKG (without case information).
+Table 6: Comparing proposed method with other RAG method
 Model Accuracy
 LLM (baseline) 0.549
+TFIDF-RAG (baseline) 0.597
+Graph-RAG (baseline) 0.179
+Light-RAG(baseline) 0.636
+LLM+CLAKG (OURS, no case) 0.676
+As shown in the Table 6, the accuracy of the LLM without external legal data is significantly higher than that of the classic model, achieving an accuracy of 0.549.
+Graph-RAG’s extremely low result may be due to its graph construction or retrieval strategy not aligning well with our dataset, indicating challenges in directly applying that method to this legal domain.
+The proposed model (LLM+CLAKG) achieved an accuracy of 0.676, which is significantly higher than that of the TFIDF-RAG, Light-RAG, Graph-RAG method.
+This indicates that the proposed approach effectively mitigates the hallucinations in LLMs.
+We believe this is because other RAG method only matches at the text level, while the latter matches the new case with keywords at the semantic level.
+Additionally, the latter utilizes all information within the CLAKG through graph embedding techniques, allowing it to calculate the most relevant law articles from a holistic perspective.
+4.4.3 Ablation Study
+In this ablation study, we evaluate the impact of incorporating different types of information on the performance of LLMs for law article recommendation tasks.
+Specifically, we compare the performance of the baseline models with and without additional case and law article information.
+We conducted experiments using both the proprietary GPT-4 model and the open-source Llama-3.1-405b-chat model.
+The baseline results for GPT-4 showed an accuracy of 0.549, while incorporating case-specific information with CLAKG (Case Law and Statutory Knowledge Graph) improved the accuracy to 0.694.
+Similarly, for the Llama-3.1-405b-chat model, the baseline performance was 0.619, and adding reference case and statutory information boosted the accuracy to 0.786.
+These results clearly demonstrate that the inclusion of relevant information, such as reference cases, significantly enhances model performance across both proprietary and open-source models.
+This validates the robustness and versatility of our approach, showing its applicability not only to high-performing proprietary models but also to widely available open-source models.
+The detailed results are summarized in Table 7.
+Table 7: Results of different models in the ablation study.
+Model Accuracy
+LLM (GPT-4, baseline) 0.549
 LLM + CLAKG (OURS, with case) 0.694
-As shown in the Table 6, the accuracy of the LLM without external legal data is significantly higher than that of the BERT model, achieving an accuracy of 0.549.
-However, it is still lower than the proposed model (LLM+CLAKG).
-We attribute this to the hallucinations of LLM and the proposed approach effectively mitigates the impact of label imbalance in training.
-The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than that of the LLM-based method.
-This indicates that the proposed approach effectively mitigates the Huallucinations in LLMs.
-3.4.3 Analyzing the Impact of Utilizing Macro Semantics
-We utilized the LLM (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods) based on TFIDF for law article recommendation as comparison method.
-This method was compared against two method proposed in this paper: the LLM for law article recommendation based on CLAKG (with case information) and the LLM for law article recommendation based on CLAKG (no case information).
-Table 7: Model Accuracy Comparison Between LLM+TFIDF and Ours
-Model Accuracy
-LLM+TFIDF (base line) 0.595
-LLM+CLAKG (OURS, no case) 0.676
-LLM+CLAKG (OURS, with case) 0.694
-As shown in the Table 7, both the TFIDF-based law article recommendation method and the CLAKG-based law article recommendation method mitigate the hallucinations of the LLM and improve its accuracy, with the latter performing better.
-We believe this is because the former only matches at the text level, while the latter matches the new case with keywords at the semantic level.
-Additionally, the latter utilizes all information within the CLAKG through graph embedding techniques, allowing it to calculate the most relevant law articles from a holistic perspective.
+LLM (llama-3.1-405b-chat, baseline) 0.619
+LLM + CLAKG (llama-3.1-405b-chat, with case) 0.786
+This ablation study highlights the significant impact that additional legal context, such as case and law article information, can have on the performance of LLMs for law article recommendation tasks, confirming the effectiveness of our proposed methodology.
 5 Conclusion

### v3

@@
 Abstract
-Court efficiency is vital for social stability.
-However, in most countries around the world, the grassroots courts face case backlogs, with decisions relying heavily on judicial personnel’s cognitive labor, lacking intelligent tools to improve efficiency.
-To address this issue, we propose an efficient law article recommendation approach utilizing a Knowledge Graph (KG) and a Large Language Model (LLM).
-Firstly, we propose a Case-Enhanced Law Article Knowledge Graph (CLAKG) as a database to store current law statutes, historical case information, and correspondence between law articles and historical cases.
-Additionally, we introduce an automated CLAKG construction method based on LLM.
-On this basis, we propose a closed-loop law article recommendation method.
-Finally, through a series of experiments using judgment documents from the website ”China Judgements Online”, we have improved the accuracy of law article recommendation in cases from 0.549 to 0.694, demonstrating that our proposed method significantly outperforms baseline approaches.
+Judicial efficiency is critical to social stability.
+However, in many countries worldwide, grassroots courts face substantial case backlogs, and judicial decisions remain heavily dependent on judges’ cognitive efforts, with insufficient intelligent tools to enhance efficiency.
+To address this issue, we propose a highly efficient law article recommendation approach combining a Knowledge Graph (KG) and a Large Language Model (LLM).
+First, we construct a Case-Enhanced Law Article Knowledge Graph (CLAKG) to store current law articles, historical case information, and their interconnections, alongside an LLM-based automated construction method.
+Building on this, we propose a closed-loop law article recommendation framework integrating graph embedding-based retrieval and KG-grounded LLM reasoning.
+Experiments on judgment documents from China Judgments Online demonstrate that our method boosts law article recommendation accuracy from 0.549 to 0.694, outperforming strong baselines significantly.
+To support reproducibility and future research, all source code and processed datasets are publicly available on GitHub (see Data Availability Statement).
 keywords:
@@
 1 Introduction
-In a modern rule-of-law society, the efficiency of court judgments is crucial for ensuring economic growth and social stability[1, 2, 3].
-Meanwhile, [4] revealed that the current backlog of cases in courts is enormous, which means the workload of judicial personnel is heavy.
-[5] showed that there is a great need for intelligent tools to assist judicial personnel in improving the efficiency of case adjudication, since intelligent tools boost judicial efficiency by categorizing similar cases, speeding up research, and staying updated with new laws, ensuring faster and more consistent decisions.
-Law article recommendation entails predicting the relevant law articles applicable to a case based on its factual description.
-With the advancement of computer science, many data science techniques have been applied to this task, greatly improving efficiency.
-Classic law article recommendation techniques, such as LSTM[6], Text-CNN[7], GRU[8], RCNN[9], HAN[10], and DPCNN[11], are based on traditional machine learning methods.
-These methods focus solely on the correspondence between case facts and law article ids, while overlooking the semantic information in law articles.
-Moreover, these methods are highly susceptible to the effects of insufficient data and imbalanced data labels.
-Therefore, there is a need to solve above problems.
-Some researchers utilized Large Language Models (LLMs) to complete the task of law article recommendation[12, 13, 14].
-By using a large LLM, users can not only retrieve the law article in numeric and textual form, but also investigate the rationale behind its applicability to the case by asking follow-up questions.
-Moreover, users have the capability to probe into the rationale for the non-selection of other analogous provisions.
-However, [15] found that directly using LLMs is prone to hallucinations, generating incorrect law article information.
-Therefore, it is necessary to introduce a fully accurate legal information and case knowledge base for auxiliary analysis.
-To mitigate the issue of hallucination, researchers introduced Retrieval-Augmented Generation (RAG) technology to enhance the generation capability of LLMs, by retrieving relevant information from external knowledge bases[16, 17].
-Currently, RAG technology has been applied in fields such as code generation[18], autonomous driving[19], and enterprise management[20].
-However, research on RAG methods in the domain of law article recommendation remains relatively underexplored.
-RAG is divided into two parts: retrieval and generation.
-In academic discussions, greater attention is often given to the retrieval component of RAG, whereas the generation process is predominantly handled by LLMs.
-Classic retrieval methods include TFIDF-based retrieval[21], BM25-based retrieval[22], and retrieval methods based on frozen BERT[23].
-However, the above methods mainly focus on the word level and rarely use macro semantics for matching.
-To deal with that problem, [24] proposed that the Knowledge Graph (KG), an unstructured database that uses a directed graph structure to store data, is highly suitable to serve as a database for RAG technology.
-[25] have pointed out that the integration of KGs with LLMs shows great potential for future research.
-[26] have utilized LLMs to automatically construct knowledge graphs, which are then used to enhance the generation capabilities of LLMs.
-This method performs well in terms of comprehensiveness and diversity of answers.
-However, the KGs generated using this method often have complex structures and lack a fixed schema, making them difficult for users to modify, such as adding new cases or removing outdated ones.
-To address the above issues, we introduce knowledge graph technology and large language model technology to improve the output accuracy in law article recommendation tasks.
-Initially, we pre-design the schema of a Case-Enhanced Law Article Knowledge Graph (CLAKG) capable of simultaneously storing current law articles and historical cases information as a database.
-On this basis, we also introduce an automated CLAKG construction method based on LLM.
-Furthermore, we design a closed-loop human-machine collaboration method to complete the law article recommendation task.
-To demonstrate the effectiveness of our proposed method, we constructed a dataset of Chinese criminal law judgments.
-We compared the method introduced in this paper with approaches based on BERT[23], GRU[8], DPCNN[11], and several RAG approaches such as TFIDF-RAG[21], Graph-RAG[27] and Light-RAG[28].
-Finally, we conducted an ablation study of our method.
+In a modern rule-of-law society, judicial efficiency is critical to safeguarding economic growth and social stability[12, 25, 17].
+Meanwhile, [21] noted that courts currently face a massive case backlog, resulting in a heavy workload for judicial personnel.
+[43] emphasized an urgent need for intelligent tools to assist judicial personnel in enhancing case adjudication efficiency, as they boost judicial efficiency by categorizing similar cases, accelerating legal research, and keeping abreast of new laws—thus ensuring faster and more consistent decisions.
+Law article recommendation involves predicting the relevant law articles applicable to a case based on its factual description.
+With the development of computer science, numerous data science techniques have been applied to this task, significantly enhancing efficiency.
+Classic law article recommendation techniques—including LSTM[22], Text-CNN[26], GRU[8], RCNN[28], HAN[46], and DPCNN[24]—are based on traditional machine learning methods.
+These methods focus exclusively on the correspondence between case facts and law article IDs, while neglecting the semantic information within law articles.
+Furthermore, these methods are highly vulnerable to insufficient data and imbalanced data labels.
+Thus, there is an urgent need to address these issues.
+Several researchers have employed Large Language Models (LLMs) for law article recommendation[45, 30, 37].
+With large LLMs, users can not only retrieve law articles in numerical and textual forms but also explore the rationale for their applicability to cases through follow-up questions.
+Furthermore, users can probe the rationale for not selecting other analogous law articles.
+However, [10] found that direct use of LLMs is prone to hallucinations, producing inaccurate law article information.
+Thus, it is essential to incorporate a fully accurate legal information and case knowledge base for auxiliary analysis.
+Beyond improving the underlying knowledge base, a parallel line of work shows that prompt engineering can itself be a powerful lever for steering LLM behaviour and mitigating hallucinations.
+Recent surveys on hallucinations in LLMs highlight that a substantial portion of erroneous outputs is prompt-sensitive and can be reduced by disciplined prompting strategies, such as explicit uncertainty instructions, decomposition prompts, and verification-style prompts [23, 13].
+Within this line of work, chain-of-thought prompting and its self-consistency variant encourage models to externalize intermediate reasoning and aggregate over multiple reasoning paths, which generally leads to more reliable answers on complex reasoning and knowledge-intensive tasks [42, 41].
+Building on the same intuition that exposing and checking intermediate reasoning improves faithfulness, verification-oriented prompting schemes such as Chain-of-Verification [15] ask the model to first draft an answer, then generate targeted follow-up questions, answer them independently, and finally revise the initial answer based on the verified evidence, thereby significantly decreasing hallucination rates in open-domain question answering.
+Complementary techniques like SelfCheckGPT [31] further exploit prompt-based self-correction by sampling alternative completions and detecting contradictions as a black-box mechanism for filtering hallucinated content.
+Taken together, these studies suggest that prompt engineering provides an essential, model-agnostic axis for hallucination mitigation that naturally complements both architectural advances and RAG-style grounding.
+Motivated by this perspective, we design CLAKG-grounded prompts with explicit role specifications and strict output constraints, so that the LLM is systematically nudged to rely on verifiable legal evidence rather than fabricating non-existent law articles.
+To mitigate hallucination, researchers have introduced Retrieval-Augmented Generation (RAG) technology to enhance LLMs’ generation capabilities by retrieving relevant information from external knowledge bases[29, 4].
+Currently, RAG technology has been applied in domains such as code generation[18], autonomous driving[11], and enterprise management[32].
+However, research on RAG-based law article recommendation remains relatively underexplored.
+RAG consists of two components: retrieval and generation.
+In academic discourse, greater attention is often paid to RAG’s retrieval component, while the generation process is primarily handled by LLMs.
+Classic retrieval methods include TF-IDF-based retrieval[1], BM25-based retrieval[35], and frozen BERT-based retrieval[3].
+However, these methods primarily focus on the lexical level and rarely leverage macro-semantic information for matching.
+To address this issue, [7] proposed that Knowledge Graphs (KGs)—structured databases that use directed graph structures to store data—are highly suitable as knowledge bases for RAG technology.
+[34] noted that integrating KGs with LLMs holds significant potential for future research.
+[16] employed LLMs to automatically construct KGs, which are then used to enhance LLMs’ generation capabilities.
+This approach achieves strong performance in terms of answer comprehensiveness and diversity.
+However, KGs generated via this approach often have complex structures and lack a fixed schema, making them difficult to modify—for example, adding new cases or removing outdated ones.
+To address the above issues, we propose to combine knowledge graph technology with large language models to improve the accuracy and interpretability of law article recommendation.
+We first pre-design the schema of a Case-Enhanced Law Article Knowledge Graph (CLAKG) that can simultaneously store current law articles and historical case information in a task-oriented manner.
+On this basis, we develop an automated CLAKG construction pipeline driven by LLMs.
+Furthermore, we build a closed-loop human–machine collaboration mechanism in which CLAKG supports retrieval-augmented prompting for LLMs, while expert feedback is continuously injected back into CLAKG to refine its structure.
+To assess the effectiveness of the proposed framework, we focus on Chinese criminal law as a representative civil-law domain and construct a dataset of criminal judgments from China Judgments Online.
+We compare our method with strong text-classification baselines based on BERT[3], GRU[8], and DPCNN[24], as well as several RAG approaches such as TFIDF-RAG[1], Graph-RAG[16], and Light-RAG[20].
+We further conduct ablation studies to investigate the contribution of different components of our framework.
+Building on the above design and empirical study, this work makes the following main contributions:
+  1. 1.
+    We propose a Case-Enhanced Law Article Knowledge Graph (CLAKG) that unifies law articles and adjudicated cases under a task-oriented schema, together with an automated CLAKG construction pipeline based on LLMs.
+  2. 2.
+    We design an LLM–CLAKG-based law article recommendation framework that integrates CLAKG with LLMs in a retrieval-augmented generation (RAG) architecture.
+By grounding the LLM in CLAKG during recommendation, the framework mitigates hallucinations.
+  3. 3.
+    We build a Chinese criminal law dataset and conduct extensive experiments showing that our method improves law article recommendation accuracy from 0.549 (LLM only) to 0.694 (LLM + CLAKG with case information), consistently outperforming strong text-classification baselines (BERT, GRU, DPCNN) and RAG baselines (TFIDF-RAG, Graph-RAG, Light-RAG).
 The remainder of this paper is organized as follows.
-Section 2 reviews related works about law article recommendation, RAG technology and knowledge graphs in legal domain.
-Section 3 proposes a closed-loop law article recommendation framework.
-Section 4 utilizes the method from the third section to construct the CLAKG and compares the accuracy of this method with other baseline approaches.
-Lastly, the conclusions and future work are highlighted in Section 5.
+Section 2 reviews related work on law article recommendation, RAG technology, and knowledge graphs in the legal domain.
+Section 3 presents the proposed closed-loop law article recommendation framework.
+Section 4 describes the construction of CLAKG and compares the accuracy of the proposed method with other baseline approaches.
+Lastly, Section 5 concludes the paper and outlines directions for future work.
 2 Related Work
@@
 Despite these gains, most prior models formulate law article recommendation as mapping facts to article IDs, without directly leveraging the content of the law articles themselves.
-This focus solely on fact-to-ID correspondence makes them heavily data-dependent and sensitive to label imbalance in training sets[33].
+This focus solely on fact-to-ID correspondence makes them heavily data-dependent and sensitive to label imbalance in training sets.
 In practice, it also limits interpretability – the models do not explain why a recommended article is relevant, since the law articles are not explicitly used in the reasoning process.
@@
 2.2 LLMs and Retrieval-Augmented Generation Techniques
-The advent of large language models (LLMs), such as GPT-style transformers, has opened new avenues for law article recommendation.
+The advent of LLMs, such as GPT-style transformers, has opened new avenues for law article recommendation.
 Rather than training a specific classifier, one can prompt an LLM with a case’s facts and ask for relevant law articles.
@@
 Legal phrasing varies, and a model might need to retrieve statutes using synonymous terms or implied connections, something neural retrieval aims to handle better than lexical methods.
-Overall, RAG represent a promising direction for law article recommendation, as they combine the reasoning ability of generative models with the grounding of information retrieval, thereby reducing hallucinations and improving the transparency of the recommendation.
+Overall, RAG methods represent a promising direction for law article recommendation, as they combine the reasoning ability of generative models with the grounding of information retrieval, thereby reducing hallucinations and improving the transparency of the recommendation.
 2.3 Knowledge Graphs in the Legal Domain
@@
 Prior works have constructed legal knowledge graphs to assist various tasks.
-For example, law-specific knowledge bases have been used to integrate charge definitions and article relationships into prediction models, yielding performance gains and better interpretability in tasks like judgment prediction[34, 35].
-More recently, Chaudhri et al.[24] argued that KGs are well-suited as the underlying database for RAG systems: instead of retrieving free text, one can query a knowledge graph to retrieve facts or linked nodes, ensuring that the generation is grounded in a consistent knowledge structure.
+For example, Chaudhri et al.[7] argued that KGs are well-suited as the underlying database for RAG systems: instead of retrieving free text, one can query a knowledge graph to retrieve facts or linked nodes, ensuring that the generation is grounded in a consistent knowledge structure.
 In the legal domain, a KG could be queried to find statutes related to certain legal concepts or past cases with similar fact patterns.
@@
 In the context of law, researchers have begun designing hybrid systems that use KGs alongside LLMs.
-Chen et al.[33], for instance, construct a case-enhanced law article KG that links statutes with relevant past cases, and then use an LLM-based pipeline to recommend articles by traversing the graph for pertinent nodes (law articles) and using the LLM to generate the final recommendations and explanations.
-Such approaches have demonstrated improved accuracy over text-only baselines, highlighting that the structured knowledge in KGs can effectively guide the retrieval and generation process.
-Meanwhile, Edge et al.[26] explore automatic construction of knowledge graphs using LLMs (generating graph triples from unstructured legal text) and using those graphs to enhance answer generation.
+Edge et al.[16] explore automatic construction of knowledge graphs using LLMs (generating graph triples from unstructured legal text) and using those graphs to enhance answer generation.
 This suggests a future where legal KGs and LLMs continuously enrich each other: the KG offers a source of truth to reduce LLM errors, and the LLM helps keep the KG up-to-date with new insights.
@@
 3 Method
-Figure 1: The pipeline of CLAKG. The Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG).
-The pipeline of CLAKG-based law article recommendation is shown in Figure 1.
+Figure 1: Overall closed-loop pipeline of the proposed CLAKG-based law article recommendation system.
+The Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Cases Knowledge Graph (ACKG).
+The overall closed-loop pipeline of the proposed CLAKG-based law article recommendation system is shown in Figure 1.
+It consists of three main stages: (i) constructing the Case-Enhanced Law Article Knowledge Graph (CLAKG) (Section 3.1); (ii) performing LLM–CLAKG based law article recommendation (Section 3.2), as shown in Figure 4; and (iii) human–machine collaboration, where legal experts review and correct the recommended articles and the CLAKG is updated accordingly (Section 3.3).
 We first establish a knowledge graph based on the designed schema.
@@
 This part is introduced in Section 3.1.
-Building upon this, we have developed a human-machine Collaborative law article recommendation framework.
-Users describe new case information, and the LLM would provide potentially applicable law articles based on keyword matching and graph embedding techniques.
-Moreover, the system can provide similar historical case information from the CLAKG, followed by providing most relevant law articles that match the new case using LLM.
-This part is introduced in Section 3.2.
-Following the above process, the user reviews the law articles provided by the system and corrects any inaccuracies.
-The new case information is used to update the CLAKG.
+Building upon this, we have developed a human–machine collaborative law article recommendation framework.
+Users describe new case information, and the LLM provides potentially applicable law articles based on keyword matching and graph embedding techniques.
+Moreover, the system can provide similar historical case information from the CLAKG, followed by providing the most relevant law articles that match the new case using LLM (Section 3.2).
+Following the above process, the legal expert reviews the law articles provided by the system and corrects any inaccuracies.
+The new case information is used to update the CLAKG, forming a closed-loop human–machine collaboration process whose practical effect is illustrated in Section 3.3.
 3.1 Case-Enhanced Law Article Knowledge Graph Construction
-Figure 2: Schema of CLAKG. Ellipses of different colors represent
-different entity types.
-Figure 3: Example of CLAKG.
+Figure 2: Schema of CLAKG. Ellipses of different colors represent different entity types.
+Figure 3: Example subgraph of CLAKG instantiated according to the schema defined in Figure 2.
 Table 1: Entity types in Case-Enhanced Law Article Knowledge Graph
@@
 Original article of law An original article from the field of law, such as criminal law or traffic law, etc
-Key information Key information in law article extracted by large language model
+Key information Key information in law article
 Law article id The article number in the law
@@
 Id A law article has a unique law article id corresponding to it
-Agree With A case agree with several pieces of key information
-Applicable Law The law mentioned in the judgment documents that matches the case
+Agree With A case agrees with several pieces of key information
+Applicable Law The law that matches the case
 Occur in Time The verdict in a case occurs on a fixed date
-Reason Case Documents contain the reason of lawsuit in the case
-Detail The name of a case corresponds to the specific information of the case
-As shown in Figure 2, the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG).
+Reason Links a case to its prosecution reason stated in the judgment
+Detail Links a case to its fact description
+Figure 2 illustrates the schema of CLAKG, defining the entity types and relation types used in our system.
+An example subgraph instantiated according to this schema is shown in Figure 3.
+Tables 1 and 2 provide the details of all types of nodes and relationships in the CLAKG.
+As shown in Figure 2, the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Cases Knowledge Graph (ACKG).
 The LAKG includes original article of law, law article id, and key information.
-The key information is extracted from the original article of law using a LLM.
+The key information is first extracted from the original article of law using an LLM, and then reviewed and, if necessary, refined by legal experts before being stored as nodes in CLAKG.
 The ACKG contains the name of a case, time of the court session, prosecution reason, and the specifics of a case.
-The description of the case process is derived from the content of the judgment documents processed using a LLM.
-Based on the reference law articles written in the judgment documents, it is possible to link the name of a case node with law article id nodes.
-We analyze the correlation between case information and the corresponding key information of applicable law articles using a LLM and ultimately select up to five of the most relevant keywords for connection.
-To ensure the consistency and accuracy of the knowledge graph, we ignore any keywords that do not exist in the graph output by the large language model.
-Tables 1 and 2 provide the details of all types of nodes and relationships in the CLAKG.
+The description of the case process is derived from the content of the judgment documents processed using an LLM, and again experts check and adjust the extracted elements so that the resulting case nodes and their attributes faithfully reflect the original judgments.
+Based on the reference law articles cited in the judgment documents, we link each name of a case node in ACKG to the corresponding law article id nodes.
+Furthermore, we analyze the correlation between the factual description of a case and the key information of the applicable law articles using an LLM, and select up to five of the most relevant key-information nodes for each case.
+Concretely, the LLM is asked to choose several items from the existing list of key-information strings for the cited articles; its answer is parsed as a semicolon-separated list and each substring is matched against this list.
+Any substring that does not exactly match an existing key-information node is discarded, and again no new nodes are added.
+After this automatic step, legal experts can confirm the suggested case–article links and key-information associations, correct inaccurate links, and add missing but legally important relations if the LLM has overlooked them.
+This human–machine collaborative filtering strategy ensures that the constructed CLAKG remains consistent with the predefined schema: noisy LLM outputs may reduce recall before expert correction, but they cannot introduce spurious concepts or inconsistent node types into the graph.
+Finally, each name of a case node is linked to several key information nodes.
 3.2 LLM and CLAKG Based Law Article Recommendation Framework
 Figure 4: The LLM and CLAKG based law article recommendation framework.
-The LLM and CLAKG based law article recommendation framework is shown in Figure 4.
+The candidate law articles and historical cases retrieved from CLAKG are used to constrain the LLM’s reasoning and thereby mitigate hallucinations.
+The LLM and CLAKG based law article recommendation framework is shown in Figure 4 and step 2 of Figure 1.
 Before the start of the law article recommendation task, the CLAKG constructed in Section 3.1 needs to undergo graph embedding preprocessing using the Relational Graph Convolutional Network (RGCN) model mentioned in Section 3.2.1.
@@
 In Section 3.2.2, a LLM is used to match the K most relevant keywords to this new case (with K=8 as selected in this paper).
-Based on this, in Section 3.2.3, a candidate law article retrieval algorithm is employed to retrieve q candidate law articles that may be suitable for the new case (with q=5 as chosen in this paper).
-In Section 3.2.4, the information from the new case input by the user is integrated with the candidate law articles and historical case information obtained in Section 3.2.3 into a prompt, which is then passed to the LLM to complete the law article recommendation task.
+Based on this, in Section 3.2.3, a candidate law article retrieval algorithm is employed to retrieve q candidate law articles that may be suitable for the new case (with q=5 as selected in this paper).
+In Section 3.2.4, the information from the new case input by the user is integrated with the candidate law articles and historical case information obtained in Section 3.2.3 into a prompt, which is then passed to the LLM to complete the law article recommendation task, outputting at most $t$ law article IDs (with $t=1$ as selected in this paper).
 3.2.1 Graph Embedding on CLAKG with RGCN
-Graph embedding involves utilizing the topological structure information in knowledge graphs to assign embedding vectors to each node within the graph.
-The Graph Convolutional Network (GCN) model, initially proposed by [36], was the first graph neural network used for graph embedding.
-[37] improved upon the GCN, introducing the Relation Graph Convolution Neural Networks (RGCN), which assigns different weights to different relationships, enhancing the quality of graph embeddings.
+Given the CLAKG constructed in Section 3.1, we now describe how to compute node embeddings on this graph.
+Let $G=(V,E,R)$ denote the CLAKG, where each node $v\in V$ belongs to one of the entity types listed in Table 1 (original article of law, law article id, key information, case name, time of the court session, prosecution reason, and case details).
+Edges $e\in E$ are labeled by relation types $r\in R$ from Table 2, including Key (linking an article to its key information), Id (linking an article to its law article id), Applicable Law (linking a case to the articles cited in the judgment), Agree With (linking a case to the key information it matches), Occur in Time (linking a case to its decision date), Reason (linking a case to the reason of lawsuit), and Detail (linking a case to its detailed fact description).
+Together, these nodes and relations encode which law articles apply to which cases and through which key information, and they constitute the structural basis on which our graph embedding module operates.
+The purpose of the graph embedding module in our framework is to exploit the heterogeneous, multi-relational structure of CLAKG and map every node $v\in V$ to a low-dimensional vector representation.
+These embeddings provide the matching capability in our framework: in Section 3.2, they are used to measure the similarity between the key-information nodes selected from a new case and all law-article nodes when ranking candidate law articles.
+To learn such representations on CLAKG, we adopt Relational Graph Convolutional Networks (RGCN) [36], which extend standard Graph Convolutional Networks (GCNs) [27] to multi-relational graphs by assigning different transformation matrices to different relation types and aggregating information from neighbors according to the graph structure.
 The node embedding vector update formula for the RGCN model is shown in Equation 1.
@@
 where $h_{i}^{(l)}$ represents the embedding vector of node $i$ in the $l$-th layer, $N_{i}^{r}$ represents all neighboring nodes of node $i$ that have relationship $r$.
-$c_{i,r}$ is a hyperparameter, which in this paper is taken as $\left|N_{i}^{r}\right|$.
+$c_{i,r}$ is a normalization constant, which in this paper is taken as $\left|N_{i}^{r}\right|$.
 Both $W_{r}^{(l)}$ and $W_{0}^{(l)}$ are parameter matrices.
@@
 Figure 5: Flowchart of RGCN training process.
-The training process of the model is shown in Figure 5.
-Firstly, the system randomly initializes the embedding vectors for all nodes and relations.
-Based on this, the system updates the node embeddings using Equation 1.
-The system then calculates the score based on the DisMult decomposition, as described in Equation 2.
-This paper selects the cross-entropy function as the loss function, as described in Equation 3.
-When the triple is a positive example, meaning the information truly exists in the knowledge graph, the value of $y$ is 1.
-When the triple is a negative example, meaning the information does not exist in the graph, the value of $y$ is 0 .
-$f(s,r,o)=e_{s}^{T}R_{r}e_{o}$ (2)
-where $e_{s}$ and $e_{o}$ represent the embedding vectors of the nodes s and o, and $R_{r}$ is the embedding vector of the relation r.
-$\operatorname{Loss}=$ $-\frac{1}{(1+\omega)|\hat{\varepsilon}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Bigg[y\log l(f(s,r,o))$ $+(1-y)\log\left(1-l(f(s,r,o))\right)\Bigg]$ (3)
-where $\mathcal{J}$ represents all the datasets, and $l$ is the logistic activation function, as described in Equation 4.
-$l(x)=\frac{1}{1+e^{-x}}=\frac{e^{x}}{e^{x}+1}$ (4)
+The training process is illustrated in Fig.
+5.
+We first randomly initialize trainable embedding vectors for all entities and relations in the knowledge graph.
+Node representations are then updated by stacking $L$ layers of relational graph convolution (RGCN) as in Eq.
+1, where each layer aggregates messages from relation-specific neighbors with appropriate normalization (e.g., by the neighborhood size) and applies a non-linear activation.
+Given the final node embeddings, we compute the plausibility score of each triple $(s,r,o)$ using the DistMult [44] factorization:
+$f(s,r,o)=\langle e_{s},e_{r},e_{o}\rangle=\sum_{k=1}^{d}e_{s,k}\,e_{r,k}\,e_{o,k},$ (2)
+where $e_{s},e_{o}\in\mathbb{R}^{d}$ denote the embeddings of the subject node $s$ and object node $o$, respectively, and $e_{r}\in\mathbb{R}^{d}$ is the embedding of relation $r$.
+We optimize the model using the binary cross-entropy (BCE) loss in Eq.
+3.
+Let $\mathcal{T}$ denote the set of training examples consisting of observed positive triples and negatively sampled (corrupted) triples generated by replacing the head or tail entity with a randomly selected entity (i.e., head/tail corruption).
+The binary label $y\in\{0,1\}$ indicates whether a triple is positive ($y=1$) or a negative sample ($y=0$):
+$\mathcal{L}=-\frac{1}{|\mathcal{T}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Big[y\log\sigma(f(s,r,o))+(1-y)\log\!\big(1-\sigma(f(s,r,o))\big)\Big],$ (3)
+where $\sigma(\cdot)$ is the logistic function defined in Eq. 4.
+$\sigma(x)=\frac{1}{1+e^{-x}}=\frac{e^{x}}{e^{x}+1}.$ (4)
 3.2.2 Matching Key Information Nodes in CLAKG with LLM
-In this section, we aim to match the relevant key information nodes within CLAKG based on current case information provided by users.
-To facilitate this process, we have designed prompts for the LLM, which include the expert role, task description, details of the newly reported case, known key information nodes, and example outputs.
-The prompts for LLM in the task of matching key information nodes in the CLAKG related to new case information is shown in Table 3.
 Table 3: Prompts for LLM in the task of matching key information nodes in the CLAKG related to new case information.
@@
 Precautions Please directly output the key information nodes, separated by semicolons (;) without including any additional content (including unnecessary punctuation marks).
-Output Example Public and private property; multiple thefts; sale; committing a crime; causing damage; seriously disrupting public order; multiple thefts; lawfully performing duties
+Output Example Public and private property; sale; committing a crime; causing damage; seriously disrupting public order; multiple thefts; lawfully performing duties
+In this section, we aim to match the relevant key information nodes within CLAKG based on current case information provided by users.
+To facilitate this process, we have designed prompts for the LLM, which include the expert role, task description, details of the newly reported case, known key information nodes, and example outputs.
+The prompt template for LLM in the task of matching key information nodes in the CLAKG related to new case information is shown in Table 3.
+During inference, we reuse the global list of key-information strings stored in CLAKG and present it to the LLM as the option set in the prompt.
+The LLM is instructed to output several items from this list, separated by semicolons.
+The output is post-processed by splitting on semicolons and performing exact matching against the list; substrings that do not match any existing key-information node are discarded and do not create new nodes.
 3.2.3 Law Article and Case Retrieval Algorithm
 Algorithm 1 Candidate Law Article Retrieval Algorithm
-1:key information node, the embedded vector dictionary obtained in Section 3.2.1
-2:5 Original articles of law node
-1:Initialize: law_article_distances = [] , law_article_set = {} $\triangleright$ Initializes a list to store distances and corresponding law articles and a set to store candidate law articles
-2:for $e\in\text{key information node}$ do
-3:  Case_set.add_node($x\mid x-[\text{key}]\rightarrow e$) $\triangleright$ $x$ is ‘Original article of penal law’ node
-4:end for
-5:for $\text{law\_article}\in\text{law\_article\_set}$ do
-6:  $\text{distance}=0$ $\triangleright$ Initializes the distance score
-7:  for $e\in\text{key information node}$ do
-8:   $\text{distance}\mathrel{+}=\operatorname{dis}(\text{emb}[\text{law\_article}],\text{emb}[\text{e}])$ $\triangleright$ Calculate the distance score
-9:    where $\operatorname{dis}(a,b)=\frac{a\cdot b}{\|a\|\times\|b\|}$ $\triangleright$ Cosine similarity formula
-10:  end for
-[…]
-12:end for
-[…]
-14:return law_article_distances[:5] $\triangleright$ Output the top five candidate law articles
+1:Query key-information nodes $K$; embedding lookup table $\mathrm{emb}[\cdot]$ (Section 3.2.1); relation type $\{\textit{Key}\}$ in CLAKG
+2:Top-q candidate $\{\textit{Original article of law}\}$ nodes (optionally with scores)
+3:Initialize: $\leftarrow\emptyset; []\mathcal{L}\leftarrow[\ ] \triangleright SS$: candidate law-article set, $\mathcal{L}$: list of (score, law_article)
+4:for $e\in K$ do
+5:  $S\leftarrow S\cup\{\,x\mid(x\xrightarrow{\{\textit{Key}\}}e)\in E\ \wedge\ x\in\{\textit{Original article of law}\}\,\}$
+6:end for
+7:for $a\in S$ do
+8:  $s\leftarrow 0$
+9:  for $e\in K$ do
+10:   $s\leftarrow s+\cos(\mathrm{emb}[a],\mathrm{emb}[e])$ $\triangleright$ $\cos(u,v)=\frac{u\cdot v}{\|u\|\|v\|}$
+11:  end for
+12:  $\mathcal{L}.\texttt{append}((s,a))$
+13:end for
+14:sort($\mathcal{L}$) by $s$ in descending order
+15:return first q law articles in $\mathcal{L}$ (or their IDs)
 In this section, we propose an algorithm for retrieving historical case information and candidate law articles by leveraging the aforementioned data to identify the law article most relevant to the current case.
-The proposed method for retrieving candidate law article involves traversing all ‘key information’ nodes obtained in 3.2.2 and retrieving the ‘Original article of law’ nodes that are connected to the ‘key information’ nodes via ‘key’ relationships.
+The proposed method for retrieving candidate law article involves traversing all ‘Key information’ nodes obtained in Section 3.2.2 and retrieving the ‘Original Article of Law’ nodes that are connected to the ‘Key Information’ nodes via ‘Key’ relationships.
 These nodes are then aggregated into a set container, with all ‘Original article of law’ nodes in the set representing the candidate law article.
 Given that the number of candidate law articles in the set container may be substantial, it is essential to select those law articles most analogous to the newly reported case for further analysis.
-To this end, each node in the container is evaluated by calculating the distance scores (cosine similarity) between the ‘Original article of penal law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 3.2.1.
-These distance scores are summed to yield a cumulative distance score between each ‘Original article of law’ node and the newly reported case.
-Ultimately, the top five ‘Original article of law’ nodes with the highest cumulative distance scores are selected as the candidate law articles, serving as reference points for the LLM in completing the law article recommendation task.
+To this end, each node in the container is evaluated by calculating the similarity scores (cosine similarity) between the ‘Original article of law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 3.2.1.
+These similarity scores are summed to yield a cumulative similarity score between each ‘Original Article of Law’ node and the newly reported case.
+Ultimately, the top five ‘Original Article of Law’ nodes with the highest cumulative similarity scores are selected as the candidate law articles, serving as reference points for the LLM in completing the law article recommendation task.
 The pseudo-code is shown in Algorithm 1.
@@
 3.2.4 LLM-Enhanced Law Article Recommendation
-Our method intends to incorporate reference case information and the original text of candidate law articles to eliminate the hallucinations of LLMs.
-After retrieving candidate law articles and relevant historical cases with the method described in Section 3.2.3, The knowledge graph is then queried for key details, and this integrated information helps determine the applicability of the candidate articles.
-The prompt used for the LLM in this chapter consists of expert roles, task descriptions, information on new criminal cases, reference case information, candidate law articles, and output examples.
-After the user describes the case information to the system, the system returns the results generated by the LLM. The user can further inquire about the results provided by the LLM.
+As illustrated in Figure 6, once the CLAKG-based retrieval stages in Sections 3.2.2 and 3.2.3 have been completed, the system has obtained three types of structured evidence for a new case: (i) the textual description of the new case provided by the user, (ii) a small set of candidate law articles retrieved from CLAKG, and (iii) a set of reference historical cases linked to these candidate law articles.
+In this subsection we describe how this information is integrated into a Large Language Model so as to perform law article recommendation while mitigating hallucinations.
+Figure 6: Flowchart of LLM-enhanced law article recommendation grounded in CLAKG.
+The LLM receives KG-retrieved candidate law articles and linked historical cases as verifiable context, which constrains its reasoning and helps mitigate LLM hallucinations.
+To explicitly leverage the retrieved candidate law articles and reference historical cases as constraints for mitigating hallucinations, we design a structured prompt for the LLM-enhanced recommendation stage.
+For each new case, the prompt specifies the expert role, task description, the factual description of the case, the retrieved candidate law articles and their matched key-information nodes, as well as the reference historical cases linked to these articles.
+In addition, we impose strict output constraints that restrict the model to selecting from the candidate set and require it to explicitly state that no suitable law article is found rather than fabricating new ones.
+The detailed prompt components for CLAKG-grounded law article recommendation are summarized in Table 4.
+Table 4: Prompts for LLM in the task of CLAKG-grounded law article recommendation.
+Component Details
+Expert Role Criminal-law expert specializing in applying Chinese criminal law articles to concrete cases.
+Task Description Given the new case description, the candidate law articles, and the reference historical cases, recommend the law articles that are most applicable to the case.
+You must choose only from the candidate law articles and must not fabricate non-existent article numbers or law articles.
+New Case Description Factual description of the new case provided by the user.
+Candidate Law Articles For each candidate article: its ID and original article of law.
+Reference Historical Cases Concise summaries of historical cases linked to each candidate article in CLAKG (the specifics of the case, prosecution reason, and judgment result).
+Output Constraints Output at most $t$ law article IDs from the candidate set, each with a brief justification.
+If none of the candidates is suitable, explicitly state that no applicable law article is found, rather than inventing a new one.
+Given this prompt, the LLM compares the new case with each candidate law article and its linked historical cases, and then produces a ranked list of recommended law articles together with short textual explanations.
+The law articles remaining after these checks are returned to the user as the Final Recommended Law Articles in Figure 6.
+The user can review these recommendations, request further explanations from the LLM (e.g., why a particular article applies or why similar articles are not chosen).
+3.3 Closed-loop Human–Machine Collaboration and CLAKG Renewal
+Algorithm 2 CLAKG Update with Human Feedback
+1:CLAKG $G=(V,E,R)$, new case $c$, system law-article recommendations $A^{\text{sys}}$, expert-validated law articles $A^{\text{exp}}$, flag accepted
+2:Updated CLAKG $G^{\prime}=(V^{\prime},E^{\prime},R)$
+[…]
+4:if $c\notin V^{\prime}$ then add $c$ to $V^{\prime}$
+5:end if
+6:add / update case-detail nodes for $c$ and link them to $c$
+7:obtain $K^{\text{sys}}(c)$ via the LLM-based keyword matching in Section 3.2.2 and expert-validated $K^{\text{exp}}(c)$
+8:if accepted = True then $\triangleright$ recommendation accepted
+9:  $E^{\prime}\leftarrow E^{\prime}\cup\{(c,a,\texttt{Applicable Law})\mid a\in A^{\text{exp}}\}\cup\{(c,k,\texttt{Agree With})\mid k\in K^{\text{exp}}(c)\}$
+[…]
+11:  $E^{\prime}\leftarrow E^{\prime}\setminus\{(c,a,\texttt{Applicable Law})\mid a\in A^{\text{sys}}\}\setminus\{(c,k,\texttt{Agree With})\mid k\in K^{\text{sys}}(c)\}$
+12:  $E^{\prime}\leftarrow E^{\prime}\cup\{(c,a,\texttt{Applicable Law})\mid a\in A^{\text{exp}}\}\cup\{(c,k,\texttt{Agree With})\mid k\in K^{\text{exp}}(c)\}$
+13:end if
+14:return $G^{\prime}=(V^{\prime},E^{\prime},R)$
+In practical deployment, the LLM–CLAKG framework operates in a closed-loop manner in collaboration with legal experts, as shown in step 3 of Figure 1 and formalized in Algorithm 2.
+After the system has produced a recommended law article and a natural-language explanation for a newly emerged case (Section 3.2), experts review the proposed result.
+This review focuses on both the correctness of the recommended article and the adequacy of the supporting reasoning.
+In parallel, the system also outputs a small set of key-information nodes (“keywords”) that it deems most relevant to the new case, obtained by reusing the LLM-based matching procedure in Section 3.2.2.
+Experts inspect and, if necessary, adjust this keyword set as well, so that both the applicable law articles and their supporting key information are explicitly validated.
+If the recommendation is confirmed to be appropriate, the corresponding case–law article pair is treated as a new, high-quality supervision signal for the knowledge graph.
+Concretely, the case node of the current judgment is added to CLAKG, and one or more “Applicable Law” edges are created between this case node and the validated law-article nodes.
+In addition, the occurrence time, prosecution reason, and a concise description of the case facts stored as case-detail nodes, are linked to the new case node via “Occur in Time”, “Reason” and “Detail” relationship.
+Building on the expert-validated keyword set, the system then connects the new case to the corresponding key-information nodes by adding “Agree With” edges, again strictly reusing only key-information nodes that already exist in CLAKG.
+These operations enrich CLAKG with up-to-date case–article and case–keyword links without modifying the underlying LLM or the retrieval algorithm, corresponding to the “accepted” branch in Algorithm 2.
+Conversely, if experts deem the recommendation partially or completely incorrect, they can either select the correct articles from the candidate list or manually specify the appropriate law articles.
+The system then updates CLAKG accordingly: erroneous “Applicable Law” edges can be removed, while new edges are established between the case node and the expert-validated law articles.
+At the same time, any misleading keyword associations proposed by the system are pruned by removing the corresponding “Agree With” edges, and new “Agree With” edges are created between the case node and the expert-corrected key-information nodes.
+This correction process is captured by the “corrected” branch in Algorithm 2, where expert feedback directly refines both the case–law article and case–keyword structure of the graph.
+This human–machine collaboration enables CLAKG to continuously absorb new cases, updated keyword associations, and law article changes, leading to a denser and more accurate graph.
+Since subsequent recommendations always start from CLAKG-based retrieval and KG-grounded prompting, an improved graph directly translates into better keyword matching in Section 3.2.2, more precise candidate law-article retrieval in Section 3.2.3, and stronger factual grounding for the LLM in Section 3.2.4.
+As a result, the closed-loop renewal of CLAKG—covering both law-article and key-information links for new cases—gradually improves recommendation quality and helps mitigate hallucinations in future LLM outputs.
 4 Experiment
 4.1 Datasets
+While the general architecture of our framework is designed to be not limited to a specific jurisdiction in principle, in this study we instantiate and evaluate it on Chinese criminal law as a concrete case study.
+Specifically, the law-article dataset employed in this section is derived from the Amendment (XI) to the Criminal Law of the People’s Republic of China [39], ratified during the 24th session of the Standing Committee of the 13th National People’s Congress.
+This amendment is systematically structured into two sections: the General Provisions and the Specific Provisions, encompassing a total of 452 articles—101 in the General law articles and 351 in the Specific law articles.
+In other words, our law article set already covers the entire universe of criminal law articles under Amendment (XI), and the number of law-article nodes in CLAKG directly reflects the full statutory space relevant to this domain; it cannot be meaningfully “expanded” without moving beyond Chinese criminal law itself.
+The CLAKG was constructed utilizing the methodology in Section 3.1.
+We discuss in Section 5 the challenges and potential of adapting this framework to other legal systems (e.g., common-law jurisdictions) and languages.
 Figure 7: The name of a case, time of the court session, prosecution reason on the China Judgments Online.
-The method proposed in this paper can be applied to any type of law articles in any country.
-For research convenience, the law articles datasets employed in this section is derived from the Amendment (XI) to the Criminal Law of the People’s Republic of China[38], ratified during the 24th session of the Standing Committee of the 13th National People’s Congress.
-This amendment is systematically structured into two sections: the General Provisions and the Specific Provisions, encompassing a total of 452 articles—101 in the General law articles and 351 in the Specific law articles.
-The CLAKG was constructed utilizing the methodology in Section 3.1.
 The criminal case data analyzed in this study were obtained from the China Judgments Online [40] platform.
 Figure 7 presents one of the cases.
-In light of the fact that the Amendment (XI) to the Criminal Law of the People’s Republic of China came into force on March 1, 2021, all judicial decisions used herein pertain to cases adjudicated after this date.
-During this period, due to concerns over information security and privacy protection, the number of publicly available criminal case judgments was limited.
-As a result, certain law articles corresponded with only a limited number of cases, while some provisions had no corresponding cases at all.
-We curated a dataset comprising 1,170 case records, which collectively address law articles from 85 articles within the Specific Provisions.
-To optimize the dataset, we divided it in a manner that maximized the diversity of law articles represented in the test set.
-Specifically, 997 case records were designated for the training set and 173 for the test set.
-The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zoro shot number = 36).
-The Adjudicated Case Knowledge Graph was subsequently constructed based on the training data, following the approach outlined in Section 3.1.
+In light of the fact that the Amendment (XI) to the Criminal Law came into force on March 1, 2021, all judicial decisions used herein pertain to cases adjudicated after this date.
+During this period, due to concerns over information security and privacy protection, the number of publicly available criminal case judgments was limited, and the publication policy is outside the control of the authors.
+As a consequence, certain law articles correspond to only a very small number of cases, while some provisions have no publicly accessible cases at all.
+To mitigate this constraint, we exhaustively collected all publicly accessible criminal judgments under Amendment (XI) that matched our statutory subset, yielding a dataset of 1,170 case records covering 85 distinct specific-law articles.
+Among them, 997 case records were designated for the training set and 173 for the test set.
+The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zero-shot number = 36).
+Despite this moderate cardinality in absolute terms, the resulting law article recommendation task is already highly non-trivial: the label space is large and severely imbalanced, and, as shown in Section 4.4, several classical text-classification baselines tend to collapse to the majority label.
+Throughout all experiments, the retriever in our framework still operates over all 452 law-article nodes in CLAKG, so retrieving the correct article is not artificially simplified by a small candidate pool.
+The Adjudicated Cases Knowledge Graph was subsequently constructed based on the training data, following the approach outlined in Section 3.1.
 4.2 CLAKG Construction and Graph Embedding
@@
 Key information 3405 Id 452
-Law article id 452 Agree with 997
+Law article id 452 Agree With 4820
 The name of a case 997 Applicable Law 4118
@@
 The result produced by the LLM is consistent with the outcome recorded in the court judgment.
-4.4 Test Result
-4.4.1 Analyzing the Impact of Label Imbalance in Training
-We utilized the BERT[23], GRU[8] and DPCNN[11] model for law article recommendation as comparison methods.
-These methods were compared against the method proposed in this paper: the LLM(specifically OpenAI’s ChatGPT-4.0) for law article recommendation based on CLAKG (with case information).
-Table 5: Model accuracy comparison between classical models and ours
+4.4 Experimental Results
+In this subsection, we organize the baselines into three comparison groups to evaluate the proposed LLM+CLAKG framework on the law article recommendation task.
+Group (i) consists of classical text-classification models that directly map judgment texts to law-article labels (e.g., BERT, GRU, and DPCNN).
+Group (ii) includes LLM-based and generic RAG-style pipelines adapted from recent retrieval-augmented generation paradigms (e.g., plain LLM prompting, TFIDF-RAG, Graph-RAG, and Light-RAG), which represent the current mainstream approach for knowledge-intensive NLP tasks.
+Group (iii) presents an LLM-agnostic ablation across backbone models, where we instantiate the same CLAKG-RAG pipeline with two representative backbones—one mainstream proprietary model (GPT-4 [33]) and one strong open-source model (Llama-3.1-405B-chat [19])—and evaluate each backbone with and without CLAKG-based augmentation.
+This three-group organization enables a clearer positioning of our method with respect to (a) traditional neural classifiers, (b) recent generic RAG baselines, and (c) the incremental benefit of CLAKG when plugged into different LLM backbones under a unified evaluation setting.
+Our primary objective here is to quantify the performance gain brought by our knowledge-graph–based RAG framework (CLAKG-RAG) for law article recommendation.
+In principle, any LLM can serve as a baseline for this task by directly prompting it with case facts and asking it to output applicable law articles.
+Given the rapid proliferation of LLM families and the non-trivial monetary and engineering cost of large-scale evaluations, a comprehensive comparison against “all” available LLMs is practically and economically infeasible.
+We therefore adopt a representative strategy: we keep the CLAKG construction, retrieval pipeline, and prompting scheme fixed, and instantiate them with a small but diverse set of competitive backbone models to demonstrate both effectiveness and generality.
+This design also reflects the model-agnostic nature of our framework: in principle, any sufficiently capable instruction-following LLM (open-source or proprietary) can be plugged into the same CLAKG-RAG pipeline in future work.
+4.4.1 Comparison with classical text-classification baselines
+To analyze the impact of data scarcity and label imbalance on law article recommendation, we first compare our framework with three widely used neural models for text classification, namely BERT [14], GRU [8] and DPCNN [24].
+All of them treat law article recommendation as a text classification problem on the raw judgment documents.
+In this setting, our method corresponds to the LLM-based recommendation framework built on CLAKG with case information.
+Table 6: Accuracy of classical models vs. our LLM+CLAKG.
 Model Accuracy
@@
 DPCNN (baseline) 0.495
-LLM + CLAKG (OURS, no case) 0.676
-As shown in the Table 5, the accuracy of the BERT (baseline) is only 0.289.
-Upon analyzing the results provided by the BERT method, we discovered that this method classified all test cases under ”Article 133.” We believe this could be due to the small size of the training set, an imbalance in training labels[40], or an inappropriate loss function.
-Even after replacing the cross-entropy loss function with the Focal Loss function, the model still classified all test cases as ”Article 133”.
-The GRU (baseline) model achieved an accuracy of 0.427.
-GRU, or Gated Recurrent Unit, is a type of recurrent neural network (RNN) that is widely used for sequential data, and this result represents its performance on the task at hand.
-The DPCNN (baseline) model, which stands for Deep Pyramid Convolutional Neural Network, performed slightly better, with an accuracy of 0.495.
-DPCNN is a convolutional neural network architecture designed for text classification tasks, and its accuracy reflects its ability to capture hierarchical patterns in the input data.
-The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than any one of the classic methods.
-This indicates that the proposed approach effectively mitigates the impact of the insufficient data and label imbalance in training.
-4.4.2 Analyzing the Impact of Hallucinations in LLMs
-We utilized the LLM[41] (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods), TFIDF-RAG[21], Graph-RAG[27] and Light-RAG[28] for law article recommendation as comparison method.
-This method was compared against the method proposed in this paper: the LLM for law article recommendation based on CLAKG (without case information).
-Table 6: Comparing proposed method with other RAG method
+LLM + CLAKG (ours, with case) 0.694
+As shown in Table 6, the accuracy of the BERT baseline is only 0.289.
+A closer inspection of the predictions reveals that BERT tends to assign almost all test cases to the majority label “Article 133”.
+This collapse is likely caused by the small size of the training set, the severe label imbalance, and the difficulty of capturing fine-grained legal distinctions from limited data.
+The GRU baseline achieves an accuracy of 0.427.
+As a recurrent model designed for sequential data, GRU can capture some contextual information in the judgment texts, but it still struggles to distinguish between articles under highly imbalanced labels.
+The DPCNN baseline, a deep pyramid convolutional architecture tailored for text classification, further improves the accuracy to 0.495 by modeling hierarchical local patterns, yet it remains far from satisfactory.
+These observations indicate that, despite the moderate dataset size, the law article recommendation task on our corpus remains challenging: the label space is large (85 distinct articles in the test set) and highly imbalanced, and classical text-classification models tend to collapse to the majority label under such conditions.
+In contrast, our LLM+CLAKG model (with case information) reaches an accuracy of 0.694, significantly outperforming all classical baselines.
+This suggests that (i) explicitly introducing legal structure via CLAKG and (ii) leveraging LLMs to reason over graph-grounded evidence can effectively mitigate the negative impact of insufficient data and label imbalance, whereas pure text-classification models are prone to majority-label bias on this dataset.
+The classical text-classification models have the advantage of being conceptually simple, computationally efficient, and easy to deploy once a sufficiently large and well-balanced labeled dataset is available.
+However, they operate purely at the document-text level and do not encode the structured relations between cases, key legal elements, and law articles.
+As a result, they are highly sensitive to label imbalance and data sparsity, and tend to collapse to majority labels in our setting.
+In contrast, our LLM+CLAKG framework explicitly injects legal structure through the knowledge graph and leverages LLM reasoning over graph-grounded evidence.
+This design yields better robustness under limited and imbalanced data, at the cost of a more complex pipeline that requires KG construction and LLM inference.
+4.4.2 Comparison with RAG-based baselines
+To investigate the impact of hallucinations in LLMs and to compare with more recent law article recommendation paradigms, we further evaluate several RAG-based methods: a plain LLM [33] (OpenAI’s ChatGPT-4.0), TFIDF-RAG [1], Graph-RAG [16] and Light-RAG [20].
+These methods instantiate retrieval-augmented generation pipelines that have recently attracted considerable attention, and here they are adapted to the law article recommendation setting.
+In this group of experiments, our method corresponds to the LLM-based recommendation framework built on CLAKG without using case nodes (only statutory information), denoted as LLM+CLAKG (ours, no case).
+Table 7: Accuracy comparison with LLM and RAG-based baselines.
 Model Accuracy
@@
 LLM + CLAKG (ours, no case) 0.676
-As shown in the Table 6, the accuracy of the LLM without external legal data is significantly higher than that of the classic model, achieving an accuracy of 0.549.
-Graph-RAG’s extremely low result may be due to its graph construction or retrieval strategy not aligning well with our dataset, indicating challenges in directly applying that method to this legal domain.
-The proposed model (LLM+CLAKG) achieved an accuracy of 0.676, which is significantly higher than that of the TFIDF-RAG, Light-RAG, Graph-RAG method.
-This indicates that the proposed approach effectively mitigates the hallucinations in LLMs.
-We believe this is because other RAG method only matches at the text level, while the latter matches the new case with keywords at the semantic level.
-Additionally, the latter utilizes all information within the CLAKG through graph embedding techniques, allowing it to calculate the most relevant law articles from a holistic perspective.
+As shown in Table 7, a plain LLM without external legal data already achieves an accuracy of 0.549, which is much higher than that of the classical text-classification models.
+This confirms that large pretrained language models are able to capture a considerable amount of legal knowledge from their pretraining corpora.
+However, the LLM still suffers from hallucinations, occasionally citing inapplicable or even nonexistent articles.
+TFIDF-RAG improves the accuracy to 0.597 by retrieving judgment documents and law texts based on lexical similarity and feeding them as additional context to the LLM.
+Light-RAG further increases the accuracy to 0.636 by using a more efficient retrieval and organization strategy.
+Both methods demonstrate the benefit of grounding the LLM on retrieved legal documents.
+In contrast, Graph-RAG performs poorly on our dataset (0.179), which we attribute to its generic graph construction and retrieval strategy not aligning well with the fine-grained structure of Chinese criminal law and the limited size of our corpus.
+Our LLM+CLAKG (ours, no case) achieves an accuracy of 0.676, outperforming the plain LLM, TFIDF-RAG, Graph-RAG and Light-RAG.
+Compared with TFIDF-RAG and Light-RAG, which mainly rely on surface-level text similarity, our approach matches new cases to law articles via key information nodes in CLAKG, and uses graph embeddings (Section 3.2.1) to exploit the topological structure information of the graph.
+As a result, candidate law articles are selected based on their structured legal relevance rather than purely lexical overlap.
+Moreover, by constraining the LLM to reason over candidate articles and historical cases retrieved from CLAKG (Section 3.2.4), the system is grounded in verifiable law articles and case information, which mitigates hallucinations in article citations and explanations more effectively than generic RAG pipelines.
+In summary, the comparison with both classical classifiers and recent RAG-based methods shows that our LLM+CLAKG framework better captures the underlying legal reasoning and achieves consistently higher recommendation accuracy.
+RAG-based methods mitigate hallucinations by retrieving external texts and are relatively easy to adapt across domains.
+TFIDF-RAG and Light-RAG work well when lexical similarity is a good proxy for relevance, but they mainly exploit surface-level word overlap and do not explicitly model legal entities or relations; their performance may degrade when key legal distinctions are not reflected in simple bag-of-words similarity.
+Graph-RAG introduces a graph structure, yet its generic document graph and retrieval strategy are not tailored to the fine-grained structure of Chinese criminal law in our corpus, which limits its effectiveness.
+In contrast, our proposed LLM+CLAKG framework constructs a task-specific CLAKG schema centered on law article recommendation.
+This schema leverages legally meaningful key information in law articles together with adjudicated case information to precisely characterize the relations among different law articles as well as the links between cases and law articles.
+During the retrieval-augmentation stage, this semantic and structured information is injected into the LLM through prompt engineering.
+In this way, we substantially mitigate hallucinations of large language models and significantly improve the accuracy of law article recommendation.
+The trade-off is that our approach requires schema design and KG construction effort, as well as maintaining the KG over time, but it delivers consistently higher accuracy of legal reasoning than these more generic RAG pipelines.
 4.4.3 Ablation Study
@@
 LLM (GPT-4, baseline) 0.549
-LLM + CLAKG (OURS, with case) 0.694
+LLM + CLAKG (GPT-4, with case) 0.694
 LLM (llama-3.1-405b-chat, baseline) 0.619
@@
 5 Conclusion
-This paper proposes an efficient law article recommendation approach using a Case-Enhanced Law Article Knowledge Graph (CLAKG) combined with a Large Language Model (LLM).
-CLAKG integrates law articles and case information.
-It is characterized by a rich topological structure formed through usage relationships and shared key information found in judgments, which enhances the effectiveness of law article recommendation tasks.
-The proposed approach integrates LLMs with CLAKG, enabling more accurate recommendations of law articles and related cases by utilizing macro semantics to mitigate LLM hallucinations.
-The effectiveness of the proposed approach is verified through comprehensive comparisons with several baseline models on the law article recommendation task.
-This approach effectively addresses challenges such as insufficient data, imbalanced labels, and LLM hallucinations, leading to a significant improvement in the accuracy of the law article recommendation task.
-In future work, we plan to expand the data volume of CLAKG, particularly by incorporating more case data to further enhance the performance of law article recommendation.
-Additionally, we will explore further applications of the CLAKG-LLM integration to improve the efficiency of court judgments.
+This paper presents an efficient law article recommendation approach that combines a Case-Enhanced Law Article Knowledge Graph (CLAKG) with a Large Language Model (LLM).
+CLAKG integrates law articles and adjudicated cases under a task-oriented schema, with a rich topological structure formed by usage relations and shared key information extracted from judgments.
+By grounding the LLM in this graph-structured macro semantics, the proposed framework produces more accurate recommendations of law articles and related cases, while mitigating typical LLM hallucinations.
+Extensive experiments against strong text-classification baselines (BERT, GRU, DPCNN) and recent RAG-based methods (TFIDF-RAG, Graph-RAG, Light-RAG) demonstrate that our approach effectively alleviates issues such as insufficient data, label imbalance, and hallucination in LLMs, leading to substantial gains in recommendation accuracy.
+Looking ahead, we plan to expand CLAKG in both scale and scope.
+As more criminal judgments under Amendment (XI) become publicly available, we will be able to evaluate the framework on larger and more diverse corpora.
+When appropriate data access is possible, extending the evaluation to other jurisdictions will further test the scalability and generality of our approach.
+Beyond passively waiting for new real-world judgments, a complementary research direction is to explore carefully controlled data augmentation via large language models: LLMs can be used to generate hypothetical criminal case fact patterns conditioned on specific law articles, after which legal experts verify, correct, or discard these synthetic cases before they are integrated into CLAKG as new training and evaluation samples.
+Such human-validated synthetic cases have the potential to mitigate label imbalance and sparsity while preserving legal plausibility.
+It is important to emphasize that the closed-loop human–machine collaboration proposed in this paper is a deployment-level mechanism for maintaining and improving CLAKG over time, rather than a separate algorithmic baseline.
+Accordingly, our quantitative experiments in Section 4 focus on recommendation accuracy under different model variants, assuming this closed-loop mechanism is in place.
+A natural next step is to design quantitative evaluation protocols for the human feedback loop itself—for example, tracking the reduction in the proportion of recommendations that require correction, measuring expert time and effort, and analyzing how the evolving CLAKG affects downstream accuracy over longer periods.
+The current CLAKG construction still relies on LLM-based extraction of key information and case descriptions.
+Although we constrain the LLM with a predefined schema and strict post-processing, extraction errors inevitably remain and primarily manifest as missing or noisy links in the graph.
+Developing more robust and formally guaranteed information-extraction pipelines (e.g., combining supervised or rule-based extraction with LLMs) is therefore an important direction for future work.
+While our experiments focus on Chinese criminal law, a civil-law system with codified statutes, the proposed CLAKG–LLM framework is conceptually applicable to other legal systems as well.
+Adapting it to different jurisdictions, however, raises several challenges.
+First, the underlying legal structure may differ substantially: common-law systems rely more heavily on precedents and citation networks between cases, which would require extending the CLAKG schema with additional entity and relation types (e.g., precedent cases, holdings, ratios, and explicit citation links) and redesigning task-specific key information nodes.
+Second, legislative drafting conventions and document structures vary across countries (e.g., hierarchical numbering, cross-references, annotations), so the extraction–transformation–loading (ETL) pipeline and schema-design steps must be tailored to each legal corpus, as illustrated by recent work on Italian legislation [38, 9].
+Third, applying the framework to other languages calls for high-quality LLMs and tokenization tools for the target language, as well as careful adaptation of prompts and extraction templates to local legal terminology.
+Despite these challenges, the overall architecture—schema-driven KG construction, graph-embedding-based retrieval, and KG-grounded LLM reasoning—remains unchanged, suggesting that our approach can serve as a blueprint for building similar systems in other jurisdictions.
+Data Availability
+All source code, the constructed Case-Enhanced Law Article Knowledge Graph (CLAKG), and the processed law article and case datasets used in this paper are publicly available in a GitHub repository maintained by the authors:
+The repository includes: (i) dump files of CLAKG, (ii) the de-identified structured data used in our experiments, and (iii) the full implementation of the proposed LLM–CLAKG law article recommendation framework together with instructions for reproducing the experimental results.
+The original raw judgments remain accessible through the official “China Judgments Online” website, subject to its access and privacy policies.